### v1

# Gradient Extrapolation for Debiased Representation Learning

Abstract
Machine learning classification models trained with empirical risk minimization (ERM) often inadvertently rely on spurious correlations. When absent in the test data, these unintended associations between non-target attributes and target labels lead to poor generalization. This paper addresses this problem from a model optimization perspective and proposes a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), designed to learn debiased representations in both known and unknown attribute training cases. GERNE uses two distinct batches with different amounts of spurious correlations to define the target gradient as the linear extrapolation of two gradients computed from each batch’s loss. It is demonstrated that the extrapolated gradient, if directed toward the gradient of the batch with fewer amount of spurious correlation, can guide the training process toward learning a debiased model. GERNE can serve as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases. The theoretical upper and lower bounds of the extrapolation factor are derived to ensure convergence. By adjusting this factor, GERNE can be adapted to maximize the Group-Balanced Accuracy (GBA) or the Worst-Group Accuracy. The proposed approach is validated on five vision and one NLP benchmarks, demonstrating competitive and often superior performance compared to state-of-the-art baseline methods.
1 Introduction
Deep learning models have demonstrated significant success in various classification tasks, but their performance is often compromised by datasets containing prevalent spurious correlations in the majority of samples [51, 29, 18, 13]. Spurious correlations refer to unintended associations between easy-to-learn non-target attributes and target labels, leading models based on Empirical Risk Minimization (ERM)— a widely used approach in classification tasks [44]— to rely on these correlations instead of the true, intrinsic features of the classes [12, 10, 40]. This occurs because the ERM objective optimizes for the average performance [44], which results in poor generalization when these spurious features are absent. For instance, in the Waterbirds classification task [45], where the goal is to classify a bird as either a waterbird or a landbird, the majority of waterbirds are associated with water backgrounds. In contrast, the majority of landbirds are associated with land backgrounds. A model trained with ERM might learn to classify the birds based on the background-water for waterbirds and land for landbirds-rather than focusing on the birds’ intrinsic characteristics. This reliance on the spurious feature allows the model to perform well on the majority training samples, where these correlations hold, but fails to generalize to test samples where these correlations are absent (e.g., waterbirds on land). Examples of Waterbirds images shown in Fig. 1$(a)$. Avoiding spurious correlations is crucial across various applications, including medical imaging [25, 37], finance [11], and climate modeling [17].
This pervasive challenge has spurred extensive research into strategies for mitigating the negative effect of spurious correlations, particularly under varying levels of spurious attribute information availability. The authors of [49] provide a comprehensive review of the methods and research directions aimed at addressing this issue. In an ideal scenario, where attribute information is available in both the training and validation sets, methods can leverage this information to counteract spurious correlations [16, 39, 47]. When attribute information is available only in the validation set, methods either incorporate this set into the training process [18, 32, 42] or restrict its use to model selection and hyperparameter tuning [31, 28, 29, 27, 35]. Despite these efforts, existing methods still struggle to fully avoid learning spurious correlations, especially when the number of samples without spurious correlations is very limited in the training dataset, leading to poor generalization on the test data where these correlations are absent.
In this paper, we adopt a different research approach, seeking to address the issue of spurious correlations from a model optimization perspective. We propose a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), to improve generalization and learn debiased representations. The contributions of this paper can be summarized as follows:
  • We propose GERNE, a novel and easy-to-implement debiasing method in classification tasks. The core idea is to sample two types of batches with varying amounts of spurious correlations and compute the two losses on these two batches. We linearly extrapolate the gradients of these two losses to obtain a target gradient. The target gradient, controlled by an extrapolation factor, is used to update the model’s parameters.
  • The proposed gradient extrapolation approach is presented theoretically as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases.
  • The extrapolation factor’s theoretical upper and lower bounds are derived to ensure convergence, and its impact on performance is experimentally discussed.
  • We also establish a link between the extrapolation factor and both the Group-Balanced Accuracy (GBA) and Worst-Group Accuracy (WGA) metrics and generalize GERNE to the case with unknown attributes.
  • We highlight that in a biased dataset, overpresenting the minority groups in the sampled batches (compared to the majority) might be beneficial and can lead to SOTA results.
  • We validate our approach on six benchmarks spanning both vision and NLP tasks, demonstrating superior performance compared to state-of-the-art methods.
Figure 1: ($a$): Sample images from the waterbirds classification task. The majority of images with waterbirds have water background (e.g., $y=2,a=2$), and the majority of images with landbirds have land background (e.g., $y=1,a=1$). This correlation between bird class and background introduces spurious correlations in the dataset. $(b)$: $B_{b}$ shows a biased batch where the majority of images from class $y=1$ (first row of the batch) have an attribute $a=1$ (yellow color) and the majority of images from class $y=2$ (second row of the batch) have an attribute $a=2$ (light-blue color). $B_{lb}$ shows a more balanced attribute distribution (controlled by $c$, where $c=\frac{1}{2}$ in this example) than $B_{b}$ within each class. $B_{rs}$ shows a group-balanced distribution and refers to batch sampled using the Resampling method[16]. $B_{ext}$ simulates GERNE’s batch when $c\cdot(\beta+1)>1$ where the minority group in the dataset is represented as majority in this batch. $(c)$: A simplified 2D representation of gradient extrapolation (i.e. $\theta\in\mathbb{R}^{2}$). $\nabla_{\theta}\mathcal{L}_{b}$ represents the gradient computed on batch $B_{b}$, and training the model with this gradient is equivalent to training with ERM objective. $\nabla_{\theta}\mathcal{L}_{lb}$ represents the gradient computed on batch $B_{lb}$. $\nabla_{\theta}\mathcal{L}_{rs}$ is the gradient computed on batch $B_{rs}$ or equivalently, an extraplated gradient for $c\cdot(\beta+1)=1$. $\nabla_{\theta}\mathcal{L}_{ext}$ is our extrapolated gradient which can be controlled by the extrapolation factor $\beta$ depending the the level of bias in the data.
2 Related Work
Debiasing according to attributes annotations availability.
Numerous studies have leveraged attribute annotations to mitigate spurious correlations and learning debiased representation [39, 54, 3, 50, 52]. For instance, Group DRO [39] optimizes model performance on the worst-case group by minimizing worst-group error during training. While these methods are effective, obtaining attribute annotations for each sample can be extremely time-consuming and labor-intensive. Consequently, recent works have explored approaches that rely on limited attribute-labeled data to reduce the dependency on full annotations [18, 42, 32]. For example, DFR [18] enhances robustness by using a small, group-balanced validation set with attribute labels to retrain the final layer of a pre-trained model. For cases where attribute information is only available for model selection and hyperparameter tuning [6, 16, 28, 31, 53], usually an initial model is used to separate samples based on the alignment between the label and spurious attributes. Samples for which this model incurs relatively low loss are considered ”easy” examples, where we expect the alignment to hold and the samples to closely resemble the majority group. In contrast, samples with high loss are considered ”hard” examples, and these samples tend to resemble the minority group [48]. This process effectively creates ”easy” and ”hard” pseudo-attributes within each class, allowing debiasing methods that traditionally rely on attribute information to be applied. For example, JTT [28] first trains a standard ERM model and then trains a second model by upweighting the misclassified training examples (hard examples) detected by the first model. Finally, a more realistic and challenging scenario arises when attribute information is entirely unavailable [4, 43]—not accessible for training, model selection, or hyperparameter tuning—requiring models to generalize without explicit guidance on non-causal features [49].
Debiasing via balancing techniques.
A prominent family of solutions to mitigate spurious correlations across the aforementioned scenarios of annotation availability involves data balancing techniques [19, 16, 46, 40, 36, 7, 21]. These methods are valued for their simplicity and adaptability, as they are typically faster to train and do not require additional hyperparameters. Resampling underrepresented groups to ensure a more balanced distribution of samples [16, 19] or modifying the loss function to adjust for imbalances [38] are common examples of these techniques.
In this work, we demonstrate in Sec. 5.3 that although balancing techniques are effective, their performance is constrained in the presence of spurious correlations. In contrast, our proposed debiasing approach mitigates the negative effects of spurious correlations by guiding the learning process in a debiasing direction, proving to be more effective.
3 Problem Setup
We consider a regular multi-class classification problem with $K$ classes and $A$ spurious attributes. Each input sample $x_{i}\in\mathcal{X}=\{x_{i}\mid i=1,\dots,N\}$ is associated with a class label $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$ and an attribute $a_{i}\in\mathcal{A}=\{1,\dots,A\}$, where $N$ is the total number of input samples in the dataset. We define a group $\mathcal{X}_{y,a}$ for $(y,a)\in\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ as the set of input samples $x_{i}$ with a class label $y$ and an attribute $a$, resulting in $|\mathcal{G}|=K\cdot A$ groups. We define $\mathcal{X}_{y}=\cup_{a\in\mathcal{A}}\mathcal{X}_{y,a}$ as the set of input samples $x_{i}$ with a class label $y$. We assume that all groups are non-empty (i.e. $\forall(y,a)\in\mathcal{G},\mathcal{X}_{y,a}\neq\emptyset$).
Our goal is to learn the intrinsic features that define the labels rather than spurious features present in a biased dataset. This ensures robust generalization when spurious correlations are absent in the test distribution. Following [39], we aim to learn a function parameterized by a neural network $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to minimize the risk for the worst-case group:
$f^{*}=\arg\min_{f}\max_{g\in\mathcal{G}}\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right].$ (1)
Where $\ell(y,f(x))\rightarrow\mathbb{R}$ is the loss function.
4 The Proposed Method: GERNE
The core idea of GERNE is to sample two batches with different amounts of spurious correlations, hereafter named the biased batch $\boldsymbol{B_{b}}$ and the less biased batch $\boldsymbol{B_{lb}}$ (Fig. 1$(b)$). Let $\mathcal{L}_{b},\mathcal{L}_{lb}$ be the losses calculated on $\boldsymbol{B_{b}}$ and $\boldsymbol{B_{lb}}$, respectively. We assume that utilizing the extrapolation of the gradients of these two losses towards the gradient of $\mathcal{L}_{lb}$ guides the model toward debiasing as illustrated in Fig. 1$(c)$. We first present our method GERNE for training with known attributes and then generalize GERNE to the unknown attribute case.
4.1 GERNE for the Known Attributes Case
In the following, we denote the joint distribution of $(y,a)$ in a sampled batch from the dataset as $p(y,a)$. We define two types of batches with different conditional attribute distributions $p(a|y)$: the biased and the less biased batches. Then, we define our target loss as the linear extrapolation of the two losses computed on the biased and less biased batches. A simplified illustration of our approach is shown in Fig. 1. Finally, we derive the link between the extrapolation factor and the risk of the worst-case group in Eq. 1, and we theoretically define the upper and lower bounds of this factor.
4.1.1 Sampling the biased and the less biased batches
The biased batch and less biased batches are sampled to satisfy the following two conditions:
  1. 1.
    Uniform sampling from classes, i.e., $\forall y\in\mathcal{Y},p(y)=\frac{1}{K}$
  2. 2.
    Uniform sampling from groups, i.e., $\forall(y,a)\in\mathcal{G},p(x|y,a)=\frac{1}{|\mathcal{X}_{y,a}|}\text{ for }x\in\mathcal{X}_{y,a}.$
The biased batch ($\boldsymbol{B_{b}}$) is sampled with a conditional attribute distribution $p_{b}(a|y)$ within each class $y$ to reflect the inherent bias present in the dataset. Specifically, $p_{b}(a|y)=\alpha_{ya}$, where:
$\alpha_{ya}=\frac{|\mathcal{X}_{y,a}|}{|\mathcal{X}_{y}|}.$ (2)
Here $|\mathcal{X}_{y,a}|,|\mathcal{X}_{y}|$ denote the cardinality of $\mathcal{X}_{y,a},\mathcal{X}_{y}$, respectively. Note that to sample a biased batch, no access to the attributes is required, and uniformly sampling from $\mathcal{X}_{y}$ for each label $y$ satisfies Eq. 2. While the less biased batch ($\boldsymbol{B_{lb}}$) is sampled with a conditional attribute distribution, denoted as $p_{lb}(a|y)$, which is defined to be less biased compared to the biased batch. That is: $\forall(y,a)\in\mathcal{G}$,
$\ \min(\frac{1}{A},p_{b}(a|y))\leq p_{lb}(a|y)\leq\max(\frac{1}{A},p_{b}(a|y)).$ (3)
Here $\mathcal{L}_{lb}$ quantifies the loss when the representation of the spurious features is partially reduced. Choosing
$p_{lb}(a|y)=(1-c)\cdot p_{b}(a|y)+c\cdot\frac{1}{A}=\alpha_{ya}+c\cdot(\frac{1}{A}-\alpha_{ya})$ (4)
satisfies the inequality in Eq. 3, where $c\in(0,1]$ is a hyperparameter that controls the degree of bias reduction. An example of the two types of batches is presented in Fig. 1$(b)$.
4.1.2 Gradient extrapolation
We define our target loss $\mathcal{L}_{ext}$ as follows:
$\mathcal{L}_{ext}=\mathcal{L}_{lb}+\beta(\mathcal{L}_{lb}-\mathcal{L}_{b}).$ (5)
Where $\beta$ is a hyperparameter, and the loss form given the joint distribution $p(x,y,a)$ is defined as:
$\mathcal{L}=\mathbb{E}_{(x,y,a)\sim p(x,y,a)}\left[\ell(y,f(x))\right].$ (6)
Given the set of parameters $\theta$ of our model $f$, the gradient vector of the loss $\mathcal{L}_{ext}$ with respect to $\theta$ can be derived from Eq. 5:
$\nabla_{\theta}\mathcal{L}_{ext}=\nabla_{\theta}\mathcal{L}_{lb}+\beta\left(\nabla_{\theta}\mathcal{L}_{lb}-\nabla_{\theta}\mathcal{L}_{b}\right).$ (7)
Our target gradient vector $\nabla_{\theta}\mathcal{L}_{ext}$ in Eq. 7 is the linear extrapolation of the two gradient vectors $\nabla_{\theta}\mathcal{L}_{lb}$ and $\nabla_{\theta}\mathcal{L}_{b}$, and accordingly, we call $\beta$ the extrapolation factor. Since the less biased batch has a less skewed conditional attribute distribution in comparison to the biased batch according to Eq. 3, the extrapolation of their gradients and toward the less biased gradient forms a new gradient ($\mathcal{L}_{ext}$) that leads to learning even more debiasing representation for some values of the extrapolation factor $\beta>0$. The concept of extraploation is shown in Fig. 1$(c)$.
4.1.3 GERNE as general framework for debiasing
Minimizing our target loss $\mathcal{L}_{ext}$ simulates minimizing the loss of class-balanced batches with the following conditional distribution of $(y,a)\in\mathcal{G}$ :
$p_{ext}(a|y)=\alpha_{ya}+c\cdot(\beta+1)\cdot\left(\frac{1}{A}-\alpha_{ya}\right).$ (8)
The proof is in Appendix A.
Based on Eq. 8, we can establish the link between GERNE and other methods for different values of $\beta$ and $c$ as illustrated in Fig. 1:
    For $\beta=-1$, $\mathcal{L}_{ext}=\mathcal{L}_{b}$ and GERNE is equivalent to balanced-classes ERM approach.
    For $c=1,\beta=0$, $p_{ext}(a|y)=\frac{1}{A}$; GERNE is equivalent to the resampling method [16] (upsampling the minority group).
    For $c\cdot(\beta+1)=1$, $p_{ext}(a|y)=\frac{1}{A}$; In this case, $\mathcal{L}_{ext}$ is equivalent, in expectation, to the loss of the resampling method ($\mathcal{L}_{rs}$ in Fig. 1). However, the variances of the losses differ. In fact, GERNE permits controlling the variance of its loss, thanks to its hyperparameters, which might help escaping sharp minima [1] and generalize better [23]. The derivation of the variance of GERNE’s loss is detailed in Appendix C.
    For $c\cdot(\beta+1)>1$, $p_{ext}(a|y)>\frac{1}{A}\text{ if }\alpha_{ya}<\frac{1}{A}$ (also $p_{ext}(a|y)<\frac{1}{A}\text{ if }\alpha_{ya}>\frac{1}{A}$). In this case, GERNE simulates batches where the underrepresented groups (i.e. $\alpha_{ya}<\frac{1}{A}$) in the dataset to be oversampled in these simulated batches.
4.1.4 Upper and lower bounds of $\beta$
Having $p_{ext}(a|y)$ in Eq. 8 in $[0,1]$, $\beta$ should satisfy:
$\begin{split}\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\min(i_{ya}^{1},i_{ya}^{2})\leq\beta\leq\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})\end{split}$ (9)
For: $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$.
These limits are used when tuning $\beta,c$.
4.1.5 Tuning $\beta$ to minimize the risk of worst-case group
Eq. 5 can be rewritten as follows:
$\mathcal{L}_{ext}=\frac{1}{K}\sum_{g=(y,a)\in\mathcal{G}}p_{ext}(a|y)(\beta)\cdot L_{g}$ (10)
where
$L_{g}=\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right].$ (11)
In the presence of spurious correlations, minority or less-represented groups often experience higher risks, primarily due to the model’s limited exposure to these groups during training [20]. Taking this into consideration, we define $g^{\prime}=(y^{\prime},a^{\prime})=\arg\min_{(y,a)\in\mathcal{G}}\alpha_{ya}$. Since $L_{g^{\prime}}$ is multiplied by $p_{ext}(a^{\prime}|y^{\prime})$ defined in Eq. 8, increasing $\beta$ assigns more weight to $L_{g^{\prime}}$ in Eq. 10 than $L_{g}$ with $g\neq g^{\prime}$, thereby encouraging a reduction in $L_{g^{\prime}}$ during training. We outline the detailed steps of our approach for the known attribute case in Algorithm 1.
Algorithm 1 GERNE for the known attribute case
Input: $\mathcal{X}_{y,a}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$ and $a\in\mathcal{A}$, $f$ with initial $\theta=\theta_{0}$, ##\## epochs $E$, batch size per label $B$, ##\## classes $K$, ##\## attributes $A$, learning rate $\eta$.
1:Choose $c\in(0,1]$, $\beta$ satisfying Eq. 9. (REC: start with $c=\frac{1}{2},\beta=1$).
2:for epoch $=1$ to $E$ do
3:     Biased Batch $B_{b}=\emptyset$, Less Biased Batch $B_{lb}=\emptyset$
4:     for $(y,a)\in\mathcal{G}$ do
5:         Sample a mini-batch $B_{b}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $\alpha_{y,a}\cdot B$;
6:         $B_{b}=B_{b}\cup B_{b}^{y,a}$
7:         Sample a mini-batch $B_{lb}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $((1-c)~{}\alpha_{y,a}+\frac{c}{A})\cdot B$
8:         $B_{lb}=B_{lb}\cup B_{lb}^{y,a}$
9:     end for
10:     Compute $\mathcal{L}_{b},\mathcal{L}_{lb}$ on $B_{b},B_{lb}$, respectively. Then, compute $\nabla_{\theta}\mathcal{L}_{b}$ and $\nabla_{\theta}\mathcal{L}_{lb}$.
11:     Compute $\nabla_{\theta}\mathcal{L}_{ext}$ using Eq. 7.
12:     Update parameters $\theta$: $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{ext}$
13:end for
4.2 GERNE for the Unknown Attributes Case
If the attributes are not known during training, it is not possible to directly sample biased and less biased batches. This is because the conditional attribute distribution, $p(a|y)$, is unknown. Hence, we follow the previous work [31, 28, 53] by training a standard ERM model $\tilde{f}$ and using its predictions to create pseudo-attributes $\tilde{a}$. By training $\tilde{f}$ using the biased batches, the model learns to rely on spurious correlations, leading to biased predictions. Leveraging these predictions, we can classify samples into hard examples (instances with low confidence) and easy examples (instances with high confidence). After training $\tilde{f}$, we select a threshold $t\in(0,1)$ and derive pseudo-attributes based on model predictions as follows: For each class $y$, we compute the predictions $\tilde{y}_{i}=p(y|x_{i})=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$. Then, we split the predictions into two non-empty subsets. The first subset contains the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ predictions, and the corresponding samples of these predictions represent group $\mathcal{X}_{y,\tilde{a}=1}$. The remaining samples forms the group $\mathcal{X}_{y,\tilde{a}=2}$. This process ensures that each set $\mathcal{X}_{y}$ is divided into two disjoint and non-empty groups. Consequently, the pseudo-attribute space consists of two values, denoted as $\mathcal{\tilde{A}}=\{1,2\}$ ($\tilde{A}=2$) with $\tilde{\mathcal{G}}=\mathcal{Y}\times\mathcal{\tilde{A}}$ replacing $\mathcal{G}$ in the unknown attribute case. $t$ is a hyperparameter and we outline the detailed steps for the unknown case in Algorithm 2 in Appendix D.
4.2.1 Tuning $\beta$ to control the unknown conditional distribution of an attribute $a$ in class $y$
After creating the pseudo-attributes and defining the pseudo-groups, suppose we form a new batch of size $B$ with $\gamma\cdot B$ samples uniformly sampled form group $(y,\tilde{a}=1)$ and $(1-\gamma)\cdot B$ samples from group $(y,\tilde{a}=2)$, for $\gamma\in[0,1]$. The conditioal attribute distribution of an attribute $a$ given $y$ in the new formed batch is given by:
$p_{B}(a|y)$ $=\sum_{\tilde{a}\in\mathcal{\tilde{A}}}p_{B}(\tilde{a}|y)\cdot p(a|\tilde{a},y)$ (12)
Because the max/min value of a linear program must occur at a vertex, we have:
$\forall\gamma\in[0,1],\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\leq p_{B}(a|y)\leq\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ (13)
This means that if: $\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)<\frac{1}{A}(\text{or }\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)>\frac{1}{A})$, then there is no value for $\gamma$ that allows creating a batch with $p_{B}(a|y)>\frac{1}{A}$ or $p_{B}(a|y)<\frac{1}{A})$ via sampling from the pseudo-groups.
Proposition 1. Creating a biased and less biased batch with the pseudo-attributes $\mathcal{\tilde{A}}=2$, and with $c,\beta$ as hyperparameter, we can simulate creating batches with more controllable conditional attribute distribution (i.e. $p_{B}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{B}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$). The proof is in Appendix E.
5 Experiments
The performance of our proposed approach, GERNE, is evaluated across five computer vision datasets and one natural language processing dataset, specifically: Colored MNIST (C-MNIST) [3, 27], Corrupted CIFAR-10 (C-CIFAR-10)[15, 31], Biased FFHQ (bFFHQ) [22, 27], Waterbird [45], CelebA [30], and CivilComments [5].
Based on the evaluation metric and experimental setup used, we divide these datasets into two groups: Datasets-1 and Datasets-2. Datasets-1 contains the first three datasets listed above, with the goal of assessing the performance of GERNE without any data augmentation methods. While Datasets-2 contains the remaining three datasets, with our implementation unified with [49]. In the following, we first describe our experimental setup and results for each group of datasets, and then we present an ablation study for hyperparameter tuning.
5.1 Experiments on Datasets-1
Datasets.
The C-MNIST dataset represents an extension of the MNIST dataset [26], incorporating colored digits. Each digit is highly correlated with a specific color, which constitutes its majority group. In the C-CIFAR-10 dataset, each category of images is corrupted with a specific type of texture noise [15]. The bFFHQ dataset comprises human face images, with ”age” and ”gender” as the target and spurious attributes, respectively. The majority of images depicting females are labeled as ”young” while the majority of images depicting males are labeled as ”old”.
Evaluation metrics.
For both C-MNIST and C-CIFAR-10, we train with varying ratios of the minority to majority examples (0.5%, 1%, 2%, and 5%), and we follow the evaluation setup of [31, 29] using GBA on the test set as our evaluation metric. For bFFHQ, we train models with a 0.5% minority ratio and evaluate the performance based on the accuracy of the minority group in line with [27].
Baselines.
We consider six baseline methods: For the known attribute case, we compare GERNE with Group DRO [39]. For the unknown attribute case, our baselines are ERM [44], JTT [28], LfF [31], DFA [27], and LC [29].
Implementation details.
We adopt the same model architectures as the baseline methods and utilize the SGD optimizer with a momentum of 0.9 and weight decay of 0.01 across all three datasets. Additional implementation details in Sec. F.1.
Results.
Tab. 1 reports our results against baseline methods for the known and unknown attribute cases. The results of baseline methods are adopted from [29]. When the attributes are known, GERNE outperforms Group DRO by a significant margin on C-MNIST and C-CIFAR-10 datasets. The improvement in performance is ranging from about 5% on C-CIFAR-10 with 5% of minority group and up to 16% on C-MNIST with 1% of minority group. Furthermore, we show that we outperform the resampling method [16] ($c=1,\beta=0$) for all ratios. More discussion added to Appendix B. For bFFHQ, GERNE results in an improvement of over 7% in comparison with the resampling method. For the unknown attribute case, our method outperforms all the baseline methods except for C-MNIST with 5% of minority group where GERNE achieves the second-best accuracy, with only a 0.18% difference from the best result, while maintaining a lower standard deviation, despite not employing any data augmentation techniques.
Table 1: Performance comparison of GERNE and baseline methods on the C-MNIST, C-CIFAR-10, and bFFHQ datasets. We report the GBA (%) and standard deviation over three trials on the test set for C-MNIST and C-CIFAR-10, with varying ratios (%) of minority samples. For bFFHQ, we report the minority group accuracy (%). Baseline results are sourced from [29] as the same experimental settings are adopted. /$\checkmark$/$\times$ / × indicate the presence/absence of attribute information in the training set, respectively. The best results are marked in bold, and the second-best are underlined.
Methods Group Info C-MNIST C-CIFAR-10 bFFHQ 0.5 1 2 5 0.5 1 2 5 Group DRO 63.12 68.78 76.30 84.20 33.44 38.30 45.81 57.32 - GERNE (ours with $c=1,\beta=0$) 77.68 $\pm$0.89 84.36 $\pm$0.21 88.15 $\pm$0.11 91.98 $\pm$0.08 45.10 $\pm$0.60 50.08 $\pm$0.42 54.85 $\pm$0.30 62.16 $\pm$0.05 69.80 $\pm$1.33 GERNE (ours) 77.79 $\pm$0.90 84.47 $\pm$0.37 88.30 $\pm$0.20 92.16 $\pm$0.10 45.34 $\pm$0.60 50.84 $\pm$0.17 55.51 $\pm$0.10 62.40 $\pm$0.27 77.10 $\pm$0.90 ERM $\times$ 35.19 $\pm$3.49 52.09 $\pm$ 2.88 65.86 $\pm$ 3.59 82.17 $\pm$ 0.74 23.08 $\pm$ 1.25 28.52 $\pm$ 0.33 30.06 $\pm$ 0.71 39.42 $\pm$ 0.64 56.70 $\pm$ 2.70 JTT $\times$ 53.03 $\pm$ 3.89 62.90 $\pm$ 3.01 74.23 $\pm$ 3.21 84.03 $\pm$ 1.10 24.73 $\pm$ 0.60 26.90 $\pm$ 0.31 33.40 $\pm$ 1.06 42.20 $\pm$ 0.31 65.30 $\pm$ 2.50 LfF $\times$ 52.50 $\pm$ 2.43 61.89 $\pm$ 4.97 71.03 $\pm$ 1.14 84.79 $\pm$ 1.09 28.57 $\pm$ 1.30 33.07 $\pm$ 0.77 39.91 $\pm$ 1.30 50.27 $\pm$ 1.56 62.20 $\pm$ 1.60 DFA $\times$ 65.22 $\pm$ 4.41 81.73 $\pm$ 2.34 84.79 $\pm$ 0.95 89.66 $\pm$ 1.09 29.75 $\pm$ 0.71 36.49 $\pm$ 1.79 41.78 $\pm$ 2.29 51.13 $\pm$ 1.28 63.90 $\pm$ 0.30 LC $\times$ 71.25 $\pm$ 3.17 82.25 $\pm$ 2.11 86.21 $\pm$ 1.02 91.16 $\pm$ 0.97 34.56 $\pm$ 0.69 37.34 $\pm$ 1.26 47.81 $\pm$ 2.00 54.55 $\pm$ 1.26 69.67 $\pm$ 1.40 GERNE (ours) $\times$ 77.25 $\pm$ 0.17 83.98 $\pm$ 0.26 87.41 $\pm$ 0.31 90.98 $\pm$ 0.13 39.90$\pm$ 0.48 45.60$\pm$ 0.23 50.19$\pm$ 0.18 56.53 $\pm$ 0.32 70.87 $\pm$ 0.61
5.2 Experiments on Datasets-2
Datasets.
We evaluate GERNE on three commonly used datasets[49]: Waterbirds [45], CelebA [30] and CivilComments [5].
Evaluation metrics.
We follow the same evaluation strategy in [49] for model selection and hyperparameter tuning: When attributes are known in both training and validation, we use the worst-group accuracy of the test set. When attributes are unknown in training, but known in validation, we use the worst-group accuracy of the validation set. When attributes are unknown in both training and validation, we use the worst-class accuracy of the validation set.
Baselines.
For each dataset, we select the three best performing methods reported in [49]. We end up with ERM[44], Group DRO[39], DFR [18], LISA [50], ReSample [19], Mixup [52], ReWeightCRT [21], ReWeight [19], CBLoss [7], BSoftmax[36] and SqrtReWeight [49]. We also report the results for CnC [53] as it adopts similar training settings.
Implementation details.
We employ the same data augmentation techniques, optimizers and pretrained models described in [49]. Further details can be found in Sec. F.2.
Results
Tab. 2 shows the worst-group accuracy of the test set for GERNE against the baseline methods under the evaluation strategy explained above. In the case of known attributes, GERNE achieves the highest performance on the CelebA and CivilComments datasets and ranks second on Waterbirds, following DFR. In case of unknown attributes in training set but known in validation, our approach again achieves the best results on the Waterbirds and CivilComments datasets and remains competitive on CelebA, closely following the top two baseline results. Notably, DFR uses the validation set to train the model, whereas GERNE only uses it for model selection and hyperparameter tuning. We include a comparison of our method’s performance against DFR, where GERNE also uses the validation set for training, in Appendix G. In the scenario where attributes are unknown in both the training and validation sets, our approach achieves the best results on the Waterbirds and CelebA datasets. However, we observe a significant drop in accuracy on CelebA compared to the second case (unknwon attributes in training but known in validation), while this drop is less pronounced on Waterbirds. This can be explained by the worst-class accuracy evaluation metric. In the validation set of CelebA, the majority examples in class 1 exhibit spurious correlations, leading to selection process to favor the majority group while disregarding the minority group. However, the validation set of Waterbirds has balanced groups within each class, resulting in only a slight performance drop for GERNE between the second and third case. This highlights the critical role of having access to the attributes in the validation set or having a group-balanced validation set for model selection and hyperparameter tuning when aiming for better results with GERNE.
Table 2: Performance comparison of GERNE and baseline methods on the Waterbirds, CelebA, and CivilComments datasets. We report the worst-group test accuracy (%) and standard deviation over three trials on the test of each dataset. Baseline results are sourced from [49] as the same experimental settings are adopted. $\checkmark$/$\checkmark$ / ✓ indicates known attributes in training and validation sets. ×$\times$/$\checkmark$ / ✓ in validation set only, and ×/$\times$/$\times$ / × in neither. Best results are highlighted in bold, and the second-best are underlined.
Methods Group Info Waterbirds CelebA CivilComments train attr./val attr. ERM ✓/✓ 69.10 $\pm$ 4.70 62.60 $\pm$ 1.50 63.70 $\pm$ 1.50 Group DRO ✓/✓ 78.60 $\pm$ 1.00 89.00 $\pm$ 0.70 70.60 $\pm$ 1.20 ReWeight ✓/✓ 86.90 $\pm$ 0.70 89.70 $\pm$ 0.20 65.30 $\pm$ 2.50 ReSample ✓/✓ 77.70 $\pm$ 1.20 87.40 $\pm$ 0.80 73.30 $\pm$ 0.50 CBLoss ✓/✓ 86.20 $\pm$ 0.30 89.40 $\pm$ 0.70 73.30 $\pm$ 0.20 DFR ✓/✓ 91.00 $\pm$ 0.30 90.40 $\pm$ 0.10 69.60 $\pm$ 0.20 LISA ✓/✓ 88.70 $\pm$ 0.60 86.50 $\pm$ 1.20 73.70 $\pm$ 0.30 GERNE (ours) ✓/✓ 90.20 $\pm$ 0.22 91.98 $\pm$ 0.15 74.65 $\pm$ 0.20 ERM $\times$/✓ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/✓ 73.10 $\pm$ 0.40 78.50 $\pm$ 1.10 69.50 $\pm$ 0.70 ReWeight $\times$/✓ 72.50 $\pm$ 0.30 81.50 $\pm$ 0.90 69.90 $\pm$ 0.60 DFR $\times$/✓ 89.00 $\pm$ 0.20 86.30 $\pm$ 0.30 63.90 $\pm$ 0.30 Mixup $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 LISA $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 BSoftmax $\times$/✓ 74.10 $\pm$ 0.90 83.30 $\pm$ 0.30 69.40 $\pm$ 1.20 ReSample $\times$/✓ 70.00 $\pm$ 1.00 82.20 $\pm$ 1.20 68.20 $\pm$ 0.70 CnC $\times$/✓ 88.50$\pm$ 0.30 88.80$\pm$ 0.90 68.90$\pm$ 2.10 GERNE (ours) $\times$/✓ 90.21 $\pm$ 0.42 86.28 $\pm$ 0.12 71.00 $\pm$ 0.33 ERM $\times$/$\times$ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/$\times$ 73.10 $\pm$ 0.40 68.30 $\pm$ 0.90 61.50 $\pm$ 1.80 DFR $\times$/$\times$ 89.00 $\pm$ 0.20 73.70 $\pm$ 0.80 64.40 $\pm$ 0.10 Mixup $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 LISA $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 ReSample $\times$/$\times$ 70.00 $\pm$ 1.00 74.10 $\pm$ 2.20 61.00 $\pm$ 0.60 ReWeightCRT $\times$/$\times$ 76.30 $\pm$ 0.20 70.70 $\pm$ 0.60 64.70 $\pm$ 0.20 SqrtReWeight $\times$/$\times$ 71.00 $\pm$ 1.40 66.90 $\pm$ 2.20 68.60 $\pm$ 1.10 CRT $\times$/$\times$ 76.30 $\pm$ 0.80 69.60 $\pm$ 0.70 67.80 $\pm$ 0.30 GERNE (ours) $\times$/$\times$ 89.88 $\pm$ 0.67 74.24 $\pm$ 2.51 63.10 $\pm$ 0.22
5.3 GERNE vs. Balancing Techniques
Balancing techniques have been shown to achieve state-of-the-art results, while remaining easy to implement [16, 49]. Notably, resampling methods often outperform reweighting strategies when combined with stochastic gradient algorithms [2]. Our results, as presented in Tab. 1, demonstrate that GERNE outperforms resampling (GERNE with $c=1,\beta=0$) when the evaluation metric is Group-Balanced Accuracy or Accuracy on minority group. More discussion in Appendix B. This highlights the flexibility of GERNE to adapt to maximize both metrics, and its superior performance in comparison to resampling and other balancing techniques, as further supported by the results in Tab. 2. In Appendix C, we provide a detailed ablation study that compares GERNE to an equivalent (in term of loss expectation)’sampling + weighting’ approach, and show how GERNE can leverage its controllable loss variance (tuned by the hyperparameters $c,\beta$) to escape sharp minima.
5.4 Ablation Study
Tuning the extrapolation factor $\beta$.
The value of $\beta$ in Eq. 7 has a significant impact on the performance of our method in debiasing the model (i.e., leading the training process in a debiased direction and avoid learning spurious features). In Fig. 2, we show how tuning $\beta$ affects the learning process in the case of the C-MNIST dataset with $0.5\%$ of minority group in the known attribute case. We show results for $\beta\in\{-1,0,1,1.2\}\text{ with }c=0.5$. For $\beta=-1$, our target loss $\mathcal{L}_{ext}$ in Eq. 5 equals the biased loss $\mathcal{L}_{b}$, which leads to learning a biased model that exhibits high accuracy in the majority group, yet demonstrates poor performance on both the minority group and the unbiased test set. As $\beta$ increases (e.g. $\beta=0,\beta=1$), the model starts learning more intrinsic features. This is evident from the improved performance on the minority group in the validation set, as well as on the unbiased test set. However, as the extrapolation factor $\beta$ continues to increase, the model begins to exhibit higher variance during the training process as shown for $\beta=1.2$, ultimately leading to divergence when $\beta$ exceeds the upper bound defined in Eq. 9 which is equal to $1.22$ in this case. By comparing the accuracies of Minority/Majority training groups in case $\beta=0,\beta=1$, we can see that both cases have around $100\%$ accuracy on minority but higher accuarcy on majority for $\beta=0$. However, we notice a better generalization when $\beta=1$. This highlights the importance of directing the training process to the right direction early in training while overfitting is expected as well.
Figure 2: The impact of tuning $\beta$ in debiasing the model $\beta\in\{-1,0,1,1.2\}$ on debiasing the model. On the left column, we plot the training losses $\mathcal{L}_{b},\mathcal{L}_{lb}$ and the target loss $\mathcal{L}_{ext}$. On the right column, we plot the average accuracy of the minority and majority groups in both training and validation, as well as the average accuracy of the unbiased test set. Each plot represents the mean and standard deviation calculated over three runs with different random seeds.
How the selection of $t$ influences the optimal value of $\beta$.
To answer this question, we conduct experiments on C-MNIST dataset with $0.5\%$ of minority group. We first train a biased model $\tilde{f}$, and use its predictions to generate the pseudo-attributes for five different values of the threshold $t$. Let’s refer to the pseudo-groups with $\tilde{a}=1$ as the pseudo-minority group. For each threshold, we tune $\beta$ to achieve the best average accuracy on test set. Simultaneously, we compute the average precision and recall for the minority group, as defined in Eq. 14. As shown in Fig. 3, with $t=0.0005$, the average precision reaches 1, indicating that all the samples in the pseudo-minority group are from the minority group. However, these samples constitute less than 20% of the total number of samples in the minority group, as indicated by the average recall. Despite this, GERNE achieves a high accuracy of approximately 70%, remaining competitive with other methods reported in Tab. 1 while using only a very limited number of minority samples ($t=$0.0005 corresponds to about 28 samples versus 249 minority samples out of 55,000 samples in the training set). As $t$ increases to 0.001 and 0.003, precision remains close to 1 while increasing the number of minority samples in the pseudo-minority group. This increase introduces more diversity among minority samples within the pseudo-minority group, allowing for lower $\beta$ values to achieve the best average accuracy on test set. However, for even higher thresholds, such as $t=$0.01, minority samples constitute less than 40% in the pseudo-minority group, prompting a need to revert to higher $\beta$ values. We conclude that identifying the minority group is of utmost importance for achieving optimal results (high average precision and high recall) and this agrees with the results presented in both Tab. 1, Tab. 2 where we achieve the best results in the case of known attributes.
Figure 3: The effect of choosing the threshold $t$ to generate the pseudo attributes on $\beta$ for best performance.
Average Precision $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ total in }(y,\tilde{a}=1)},$ (14)
Average Recall $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ minority in }(y)}$
6 Conclusion
In this work, we introduce GERNE, a novel debiasing approach designed to mitigate spurious correlations. By sampling two types of batches with varying amounts of spurious correlations, we define a debiasing direction using the gradients of the losses computed on these two batches. The extrapolated gradient vector, controlled by the extrapolation factor, is then used to update the model parameters. We demonstrate that GERNE can be adapted to optimize the Group-Balanced Accuracy (GBA) metric or the Worst-Group Accuracy (WGA) metric, leading to improved generalization and superior performance compared to state-of-the-art methods on both known and unknown attribute cases, without relying on data augmentation techniques, on multiple benchmarks. Notably, GERNE provides a general framework that encompasses common methods such as ERM and resampling, making it a valid approach for training models even on unbiased datasets. Furthermore, GERNE allows for control over the degree of loss variance during training, increasing the likelihood of escaping sharp minima and consequently improving generalization. Although effective, GERNE remains limited in its utilization of the diversity within the majority group, where spurious correlations are present. This raises a new research question: Can we identify a more effective debiasing direction based on the two previously mentioned gradients? Future work will aim to address this question. Additionally, we will explore strategies to dynamically update the extrapolation factor during training to reduce reliance on hyperparameter tuning, and investigate novel methods to better estimate the attributes in the unknown attribute case.
Supplementary Material
Appendix A Conditional attribute distribution of the extrapolated loss: Proof of Eq. 8
$\mathcal{L}$ in Eq. 6 can be written as:
[…]
Therefore, $\mathcal{L}_{b},\mathcal{L}_{lb}$ can be written as:
[…]
Where $p_{b}(a|y),p_{lb}(a|y)$ defined in Eq. 2, Eq. 4, respectively. Substituting Eq. 15, Eq. 16 in Eq. 5:
$\mathcal{L}_{ext}$ $=\int\ell(y,f(x))\,p(y)\,\Big{(}p_{lb}(a|y)$
$\quad+\beta\cdot(p_{lb}(a|y)-p_{b}(a|y))\Big{)}\cdot p(x|a,y)\,dx\,dy\,da$
$=\int\ell(y,f(x))\,p(y)\,\Big{(}\alpha_{ya}+c\cdot(\beta+1)\cdot(\frac{1}{A}-\alpha_{ya})\Big{)}\cdot$
$\quad p(x|a,y)\,dx\,dy\,da$
[…]
Then:
$p_{ext}(a|y)=\alpha_{ya}+c\cdot(\beta+1)\cdot(\frac{1}{A}-\alpha_{ya}).$
Appendix B GERNE versus the resampling method [16]
In Tab. 1, GERNE achieves higher GBA compared to the special case of GERNE with $c=1,\beta=0$ (resampling method) for both C-MNIST and C-CIFAR-10 datasets. Our explanation behind GERNE superior performance is that resampling method tend to present the majority and minority groups equally in the batch, and the model $f$ tends to prioritize learning the easy-to-learn spurious features associated with the majority group. For instance, the color in C-MNIST. While GERNE undermines learning the spurious correlations by directing the learning process more in the debiasing direction thanks to the extrapolation factor.
Appendix C GERNE versus an equivalent sampling and weighting approach
We compare GERNE with an equivalent (in term of loss expectation) sampling+weighting method, which we refer to as ’SW’. For simplicity, we assume the following:
  1. 1.
    A binary classification task where the number of classes equals the number of attributes (i.e. $K=A=2$).
  2. 2.
    The attributes are known, and the classes are balanced. (i.e. $|\mathcal{X}_{y=1}|=|\mathcal{X}_{y=2}|$).
  3. 3.
    The majority of samples which hold the spurious correlation in each class are aligned with the class label (i.e. $|\mathcal{X}_{y,a=y}|>|\mathcal{X}_{y,a\neq y}|$).
  4. 4.
    The dataset is highly biased. In other words, $\frac{|\mathcal{X}_{y,a\neq y}|}{|\mathcal{X}_{y,a=y}|}\ll 1$.
  5. 5.
    In a highly biased dataset, best performance is coupled with overpresenting the minority (conflicting samples according to assumption 1.) in early stages of training. Therefore, an overfitting on the minority is expected before the overfitting on the majority.
We refer to the expected loss of the majority (aligned) samples as $\mathcal{L}_{A}$, and the expected loss of the minority (conflicting) as $\mathcal{L}_{C}$. For GERNE, we sample two batches: biased and less biased batch, each of size $B$. From Eq. 5, $\mathcal{L}_{ext}$ can be written as:
$\mathcal{L}_{ext}=(1+\beta)\cdot\mathcal{L}_{lb}-\beta\cdot\mathcal{L}_{b}$ (17)
Since the biased batch reflects the inherent bias present in the dataset, under the third assumption, we can approximate $\mathcal{L}_{b}$ by $\mathcal{L}_{A}$, neglecting the loss on the very few conflicting samples in the batch. Therefore, we have:
$\mathcal{L}_{b}\approx\mathcal{L}_{A}$ (18)
Following the third assumption and the conditional attribute distribution in Eq. 4, we can approximate the composition of the less biased batch as follows: a proportion of $(1-\frac{c}{2})$ of the samples in the less biased batch are drawn from the aligned samples, while a proportion of $\frac{c}{2}$ of the samples from the minority group. this leads to the following approximation:
$\mathcal{L}_{lb}\approx(1-\frac{c}{2})\cdot\mathcal{L}_{A}+\frac{c}{2}\cdot\mathcal{L}_{C}$ (19)
Substituting Eq. 18, Eq. 19 into Eq. 17:
$\mathcal{L}_{ext}\approx\frac{2-c\cdot(1+\beta)}{2}\cdot\mathcal{L}_{A}+\frac{c\cdot(1+\beta)}{2}\cdot\mathcal{L}_{C}.$ (20)
We consider the following ’SW’ approach:
  • Sampling step : we sample an ’SW’ batch of size $B$ similar to the less biased batch in GERNE. Where $(1-\frac{c}{2})$ of the batch samples are from the majority group (aligned samples) and $\frac{c}{2}$ from minority (conflicting samples).
  • Weighting step: we compute the loss $\mathcal{L}_{sw}$ over the sampled batch as follows:
    $\mathcal{L}_{sw}=w\cdot\mathcal{L}_{A}+(1-w)\cdot\mathcal{L}_{C},w=\frac{2-c\cdot(1+\beta)}{2}$ (21)
    where $\mathcal{L}_{A}$ is computed over aligned samples in the ’SW’ batch and $\mathcal{L}_{C}$ is computed over the conflicting samples.
Let’s compute the variance of the two losses:
$Var(\mathcal{L}_{sw})$ $=w^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})+(1-w)^{2}\cdot Var(\mathcal{L}_{C}^{c/2})$
$\quad+2\cdot w\cdot(1-w)\cdot Cov(\mathcal{L}_{A}^{1-c/2},\mathcal{L}_{C}^{c/2})$ (22)
where $Var(\mathcal{L}^{m})$ means the variance computed over $m\cdot B$ samples where $B$ is the batch size. For simplicity, we refer to $Var(\mathcal{L}^{1})\text{ as }Var(\mathcal{L})$
Following the fourth assumption, when the model overfits on the conflicting samples (i.e. $\mathcal{L}_{C}\approx 0$), we can approximate both $Var(\mathcal{L}_{C}),Cov(\mathcal{L}_{A},\mathcal{L}_{C})$ to zero. Therefore:
$Var(\mathcal{L}_{sw})$ $\approx w^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=\frac{w^{2}}{1-\frac{c}{2}}\cdot Var(\mathcal{L}_{A})=$
$\quad(\frac{2-c\cdot(1+\beta)}{2})^{2}\cdot\frac{2}{2-c}\cdot Var(\mathcal{L}_{A})$ (23)
From Eq. 17:
$Var(\mathcal{L}_{ext})$ $=(1+\beta)^{2}\cdot Var(\mathcal{L}_{lb})+\beta^{2}\cdot Var(\mathcal{L}_{b})$
$\quad-2\cdot(1+\beta)\cdot\beta\cdot Cov(\mathcal{L}_{lb},\mathcal{L}_{b})$
$\quad\geq((1+\beta)\cdot\sqrt{Var(\mathcal{L}_{lb})}-\beta\cdot\sqrt{Var(\mathcal{L}_{b})})^{2}$ (24)
Note that the the inequality reduces to an equality in Appendix C if $Cov(\mathcal{L}_{lb},\mathcal{L}_{b})=\sqrt{Var(\mathcal{L}_{lb})}\cdot\sqrt{Var(\mathcal{L}_{b})}$.
The covariance term ($Cov$) can be controlled by the number of shared samples between the biased and less biased batches. If all the aligned samples in the less biased batch are included in the sampled biased batch (i.e. the less biased batch is created by replacing some aligned samples by conflicting ones), we get maximum value for the $Cov$.
From Eq. 19:
$Var(\mathcal{L}_{lb})\approx(1-\frac{c}{2})^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=(1-\frac{c}{2})\cdot Var(\mathcal{L}_{A})$ (25)
And from Eq. 18
$Var(\mathcal{L}_{b})\approx Var(\mathcal{L}_{A})$ (26)
Finally, substituting Eq. 25 and Eq. 26 in Appendix C:
$Var(\mathcal{L}_{ext})\geq((1+\beta)\cdot\sqrt{1-\frac{c}{2}}-\beta)^{2}\cdot Var(\mathcal{L}_{A})$ (27)
According to the fourth assumption, we are interested in the range where $c\cdot(\beta+1)\geq 1$. Using the limits of $\beta$ defined in Eq. 9, we obtain $\beta\in\left[\frac{1-c}{c},\frac{2-c}{c}\right]$. As $\beta\to\frac{2-c}{c}$, the representation of aligned samples simulates a vanishing representation (according to Eq. 8) in the sampled batches, which leads to $\mathcal{L}_{A}>0$. Assuming a limited and non-vanishing variance $Var(\mathcal{L}_{A})$ (i.e. the model outputs a non-constant prediction for samples from the majority group), we have:
$\beta\to\frac{2-c}{c}\implies Var(\mathcal{L}_{sw})\approx 0,Var(\mathcal{L}_{ext})\neq 0$ for $c\in(0,1]$. This non-vanishing variance of GERNE’s loss, if controlled with tuning $\beta$ to ensure stability, gives the model the chance of escape sharp minima similar to gradient random perturbation [1] and therefore, improve generalization [33, 23, 34].
Appendix D Algorithm 2.
Algorithm 2 GERNE for the unknown attribute case
Input: $\mathcal{X}_{y}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$, $f$ with initial $\theta=\theta_{0},\tilde{\theta}=\tilde{\theta_{0}}$ (parameters of the biased model $\tilde{f}$), ##\## epochs $E$, batch size per class label $B$, ##\## classes $K$, ##\## attributes $\tilde{A}=2$, learning rate $\eta$.
1:Training $\tilde{f}$ on biased batches with class balanced accuracy CBA $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\mathbb{P}_{x|y}(y=\arg\max_{y^{\prime}\in\mathcal{Y}}\tilde{f}_{y^{\prime}}(x))\quad$ as the evaluation metric for model selection.
2:Select a threshold $t$ and create the pseudo-groups $\tilde{\mathcal{G}}$ by following the steps in Sec. 4.2.
3:Follow Algorithm 1. with: $\mathcal{G}\leftarrow\tilde{\mathcal{G}}$.
Appendix E Proposition 1.
Creating both biased and less biased batches using the pseudo-groups $\mathcal{\tilde{G}}$, and with $\beta$ as hyperparameter, we can simulate batches with a more controllable conditional attribute distribution. Specifically, for $(y,a)\in\mathcal{G}$, we can achieve scenarios where $p_{ext}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{ext}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ as opposed to Eq. 13.
Proof. We define $\tilde{\alpha}_{y\tilde{a}}$ the same way as in Eq. 2 for the created pseudo-groups: $\tilde{\alpha}_{y\tilde{a}}=\frac{|\mathcal{X}_{y,\tilde{a}}|}{|\mathcal{X}_{y}|}$. For a constant $c$, we create the less biased batch as in Eq. 4:
$p_{lb}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (28)
Similar to Eq. 8, the conditional attribute distribution $p_{ext}(\tilde{a}|y)$ is given by:
$p_{ext}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (29)
We can write $p_{ext}(a|y)$ as follows:
$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}p_{ext}(\tilde{a}|y)\cdot p(a|\tilde{a},y).$ (30)
Placing Eq. 29 in Eq. 30, we get
$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdot p(a|\tilde{a},y)+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdot p(a|\tilde{a},y).$ (31)
For $p(a|\tilde{a}=1,y)\neq p(a|\tilde{a}=2,y)$ and $\tilde{\alpha}_{y1}\neq\frac{1}{2}$, to make $p_{ext}(a|y)=p$ for some $p\in[0,1]$, we can choose:
$\beta=\frac{p-\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdotp(a|\tilde{a},y)}{\sum_{\tilde{a}\in\tilde{\mathcal{A}}}c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdotp(a|\tilde{a},y)}-1.\blacksquare$ (32)
Discussion. When $\tilde{\alpha}_{y1}=\frac{1}{2}$, our algorithm is equivalent to sampling uniformly from $\mathcal{X}_{y}$ and equally from classes. When $p(a|\tilde{a}=1,y)=p(a|\tilde{a}=2,y)$, it implies that $\tilde{f}$ has distributed the samples with attribute $a$ and class $y$ equally between the two pseudo-groups. However, in practice, this is precisely the scenario that $\tilde{f}$ is designed to avoid. Specifically, if $a$ represents the presence of spurious attributes (i.e., the majority group), it is likely that $p(a|\tilde{a}=1,y)<p(a|\tilde{a}=2,y)$. Conversely, when $a$ represents the absence of spurious features (i.e., the minority group), we would expect $p(a|\tilde{a}=1,y)>p(a|\tilde{a}=2,y)$. In fact, $\tilde{f}$ is explicitly trained to exhibit a degree of bias, which inherently disrupts the equality above.
Appendix F Implementation Details
F.1 Implementation details on Datasets-1
For the C-MNIST [3, 27], we deploy a multi-layer perceptron (MLP) with three fully connected layers, while for C-CIFAR-10 [15, 31] and bFFHQ [22, 27], we employ a pretrained ResNet-18 model [14] (pretrained on ImageNet1K [8]) as the backbone. The Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9 and a weight decay of 0.01, is applied across all three datasets. Batch sizes are configured as follows: 100 per group/pseudo-group for C-MNIST and C-CIFAR-10, and 32 for bFFHQ. Learning rates are set to 0.1 for C-MNIST in the known attribute case and 0.01 in the unknown attribute case. For C-CIFAR-10 and bFFHQ, the learning rate is set to 0.0001.
For GERNE in the known attribute case, we present results from two experimental configurations. In the first experiment, we set $c=1\text{ and }\beta=0$, which corresponds to resampling [16] from groups, i.e., training on $\mathcal{L}_{lb}$ without extrapolation. In the second experiment, $\beta$ is tuned for $c\in\{\frac{1}{2},1\}$. In the unknown attribute case, $t$ is an additional hyperparameter to be tuned. We avoid using any data augmentations as certain transformations can unintentionally fail to preserve the original label. For example, flips and rotations in C-MNIST can distort labels (e.g., a rotated ”6” appearing as a ”9”) [41]. For training $\tilde{f}$ in case of unknown attributes in the training set, we employ the same model architecture as $f$, with modifications to the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the learning rate used to train $f$. The loss function used is the cross-entropy loss in all the experiments.
F.2 Implementation details on Datasets-2
To ensure a fair comparison of GERNE with other methods in [49], we adhere to the same experimental settings. For the Waterbirds [45] and CelebA [30] datasets, we utilize a pretrained ResNet-50 model [14] as the backbone, while for CivilComments [5], we use a pretrained BERT model [9]. Each backbone is followed by an MLP layer with $K$ output neurons. We employ SGD with a momentum of 0.9 and a weight decay of 0.01 for Waterbirds and CelebA, while for CivilComments, we use AdamW [24] optimizer with a weight decay of 0.0001 and a tunable dropout rate. We set batch sizes to 32 for both Waterbirds and CelebA and 5(16) per group(pseudo-group) for CivilComments. The learning rates are configured as follows: 0.0001 for Waterbirds and CelebA, and 0.00001 for CivilComments. Additionally, we set the bias reduction factors $c$ to 0.5 for Waterbirds and CelebA and to 1 for CivilComments. For image datasets, we resize and center-crop the images to 224×224 pixels. In the case of unknown attributes in the training set, $\tilde{f}$ has the same architecture as $f$, but we adjust the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the value used to train $f$. We employ the Cross-entropy loss as the loss function across all experiments. For experiments with unknown attributes in both the training and validation sets, we limit the hyperparameter $t$ search space to the interval $[0,\frac{1}{2}]$
Appendix G Evaluating GERNE under limited attribute information
To further demonstrate the effectiveness of GERNE in scenarios with limited access to samples with attribute information, we conduct two experiments on the CelebA dataset. In these experiments, we exclude the training set and only use the validation set with its attribute information for training. We follow the same settings and implementation details described earlier. As part of the implementation, we first tune the hyperparameters using the designated evaluation metric. Once we determine the optimal hyperparameters, we fix them and train the model $f$ three times with different random seeds. Finally, we report the average worst-group test accuracy and standard deviation across these runs.
Experiment 1 - Evaluation on test set. In this experiment, $f$ is trained using the full validation set. The worst-group accuracy on the test set is used as the evaluation metric. This setup represents the best possible performance achievable when relying solely on the validation set for training.
Experiment 2 - Cross-validation. we divide the validation set into three non-overlapping folds, ensuring that each fold preserves the same group distribution as the original validation set (i.e., we randomly and equally distribute samples from each group across the folds). We use two folds to train $f$ and reserve the remaining fold for hyperparameter tuning and model selection, using the worst-group accuracy on this fold as the evaluation metric. We repeat this process three times, with each fold serving as the validation fold exactly once. We summarize the average worst-group test accuracy and standard deviation across all nine runs (three folds × three seeds) in Tab. 3.
We also compare these results with DFR, a method that trains the last layer of the model on the validation set after performing ERM training on the training set. GERNE consistently achieves state-of-the-art results, demonstrating its robustness and effectiveness even with severely limited attribute information.
Table 3: Performance Comparison of GERNE and DFR using the validation set for training.
Method WGA on test set(%)
DFR 86.30 $\pm$ 0.30
GERNE - Evaluation on test set 90.97 $\pm$ 0.35
GERNE - Cross-validation 88.63 $\pm$ 0.59

### v2

@@
 This paper addresses this problem from a model optimization perspective and proposes a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), designed to learn debiased representations in both known and unknown attribute training cases.
-GERNE uses two distinct batches with different amounts of spurious correlations to define the target gradient as the linear extrapolation of two gradients computed from each batch’s loss.
-It is demonstrated that the extrapolated gradient, if directed toward the gradient of the batch with fewer amount of spurious correlation, can guide the training process toward learning a debiased model.
-GERNE can serve as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases.
-The theoretical upper and lower bounds of the extrapolation factor are derived to ensure convergence.
-By adjusting this factor, GERNE can be adapted to maximize the Group-Balanced Accuracy (GBA) or the Worst-Group Accuracy.
-The proposed approach is validated on five vision and one NLP benchmarks, demonstrating competitive and often superior performance compared to state-of-the-art baseline methods.
+GERNE uses two distinct batches with different amounts of spurious correlations and defines the target gradient as a linear extrapolation of the gradients computed from each batch’s loss.
+Our analysis shows that when the extrapolated gradient points toward the batch gradient with fewer spurious correlations, it effectively guides training toward learning a debiased model.
+GERNE serves as a general framework for debiasing, encompassing ERM and Resampling methods as special cases.
+We derive the theoretical upper and lower bounds of the extrapolation factor employed by GERNE.
+By tuning this factor, GERNE can adapt to maximize either Group-Balanced Accuracy (GBA) or Worst-Group Accuracy (WGA).
+We validate GERNE on five vision and one NLP benchmarks, demonstrating competitive and often superior performance compared to state-of-the-art baselines.
+The project page is available at: https://gerne-debias.github.io/.
 1 Introduction
 Deep learning models have demonstrated significant success in various classification tasks, but their performance is often compromised by datasets containing prevalent spurious correlations in the majority of samples  [52, 29, 18, 13].
-Spurious correlations refer to unintended associations between easy-to-learn non-target attributes and target labels, leading models based on Empirical Risk Minimization (ERM)— a widely used approach in classification tasks [44]— to rely on these correlations instead of the true, intrinsic features of the classes [12, 10, 40].
-This occurs because the ERM objective optimizes for the average performance [44], which results in poor generalization when these spurious features are absent.
+Spurious correlations refer to unintended associations between easy-to-learn, non-target attributes and target labels.
+These correlations often cause models trained with Empirical Risk Minimization (ERM) [45] to rely on these correlations instead of the true, intrinsic features of the classes  [12, 10, 40].
+This occurs because the ERM objective optimizes for the average performance  [45], thereby biasing the model toward the easy-to-learn features that are predictive for the majority of training samples.
+As a result, ERM-trained models often exhibit poor generalization when these spurious features are absent in the test data.
 For instance, in the Waterbirds classification task  [46], where the goal is to classify a bird as either a waterbird or a landbird, the majority of waterbirds are associated with water backgrounds.
 In contrast, the majority of landbirds are associated with land backgrounds.
-A model trained with ERM might learn to classify the birds based on the background-water for waterbirds and land for landbirds-rather than focusing on the birds’ intrinsic characteristics.
-This reliance on the spurious feature allows the model to perform well on the majority training samples, where these correlations hold, but fails to generalize to test samples where these correlations are absent (e.g., waterbirds on land).
+A model trained with ERM might learn to classify the birds based on the background—water for waterbirds and land for landbirds—rather than focusing on the birds’ intrinsic characteristics.
+This reliance on the spurious feature allows the model to perform well on the majority of training samples, where these correlations hold, but fails to generalize to test samples where these correlations are absent (e.g., waterbirds on land).
 Examples of Waterbirds images shown in Fig.
@@
 Avoiding spurious correlations is crucial across various applications, including medical imaging  [25, 37], finance  [11], and climate modeling  [17].
-This pervasive challenge has spurred extensive research into strategies for mitigating the negative effect of spurious correlations, particularly under varying levels of spurious attribute information availability.
+This pervasive challenge has spurred extensive research into strategies for mitigating the negative effects of spurious correlations, particularly under varying levels of attribute information availability.
 The authors of  [50] provide a comprehensive review of the methods and research directions aimed at addressing this issue.
@@
 In this paper, we adopt a different research approach, seeking to address the issue of spurious correlations from a model optimization perspective.
-We propose a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), to improve generalization and learn debiased representations.
+We propose a novel method, Gradient Extrapolation for Debiased Representation Learning (GERNE), to avoid reliance on spurious features and learn debiased representations.
+The core idea is to sample two types of batches with varying amounts of spurious correlations (Fig.
+1(b)) and compute the two losses on these two batches.
+We assume that the difference between the gradients of these losses captures a debiasing direction.
+Therefore, we define our target gradient as the linear extrapolation of these two gradients toward the gradient of the batch with fewer amount of spurious correlations (Fig.
+1(c)).
 The contributions of this paper can be summarized as follows:
-  • We propose GERNE, a novel and easy-to-implement debiasing method in classification tasks.
-The core idea is to sample two types of batches with varying amounts of spurious correlations and compute the two losses on these two batches.
-We linearly extrapolate the gradients of these two losses to obtain a target gradient.
-The target gradient, controlled by an extrapolation factor, is used to update the model’s parameters.
-  • The proposed gradient extrapolation approach is presented theoretically as a general framework for debiasing with methods, such as ERM, reweighting, and resampling, being shown as special cases.
-  • The extrapolation factor’s theoretical upper and lower bounds are derived to ensure convergence, and its impact on performance is experimentally discussed.
-  • We also establish a link between the extrapolation factor and both the Group-Balanced Accuracy (GBA) and Worst-Group Accuracy (WGA) metrics and generalize GERNE to the case with unknown attributes.
-  • We highlight that in a biased dataset, overpresenting the minority groups in the sampled batches (compared to the majority) might be beneficial and can lead to SOTA results.
-  • We validate our approach on six benchmarks spanning both vision and NLP tasks, demonstrating superior performance compared to state-of-the-art methods.
-Figure 1: ($a$): Sample images from the waterbirds classification task.
-The majority of images with waterbirds have water background (e.g., $y=2,a=2$), and the majority of images with landbirds have land background (e.g., $y=1,a=1$).
+  • We propose GERNE as a general framework for debiasing, with methods such as ERM and Resampling being shown as special cases.
+  • We derive the theoretical upper and lower bounds of the extrapolation factor and establish a direct connection between the extrapolation factor and the risk for the worst-case group.
+We show that tuning this factor within these bounds enables GERNE to adaptively optimize for either Group-Balanced Accuracy or Worst-Group Accuracy.
+  • We validate our approach on six benchmarks spanning both vision and NLP tasks, under both known and unknown attribute cases, demonstrating competitive and often superior performance compared to state-of-the-art methods—particularly in scenarios where samples without spurious correlations are scarce.
+Figure 1: 1(a): Sample images from the waterbirds classification task.
+Most landbird images appear with land backgrounds (i.e., $y=1,a=1$), while most waterbird images appear with water backgrounds (i.e., $y=2,a=2$).
 This correlation between bird class and background introduces spurious correlations in the dataset.
-$(b)$: $B_{b}$ shows a biased batch where the majority of images from class $y=1$ (first row of the batch) have an attribute $a=1$ (yellow color) and the majority of images from class $y=2$ (second row of the batch) have an attribute $a=2$ (light-blue color).
-$B_{lb}$ shows a more balanced attribute distribution (controlled by $c$, where $c=\frac{1}{2}$ in this example) than $B_{b}$ within each class.
-$B_{rs}$ shows a group-balanced distribution and refers to batch sampled using the Resampling method[16].
-$B_{ext}$ simulates GERNE’s batch when $c\cdot(\beta+1)>1$ where the minority group in the dataset is represented as majority in this batch.
-$(c)$: A simplified 2D representation of gradient extrapolation (i.e.
-$\theta\in\mathbb{R}^{2}$).
-$\nabla_{\theta}\mathcal{L}_{b}$ represents the gradient computed on batch $B_{b}$, and training the model with this gradient is equivalent to training with ERM objective.
-$\nabla_{\theta}\mathcal{L}_{lb}$ represents the gradient computed on batch $B_{lb}$.
-$\nabla_{\theta}\mathcal{L}_{rs}$ is the gradient computed on batch $B_{rs}$ or equivalently, an extraplated gradient for $c\cdot(\beta+1)=1$.
-$\nabla_{\theta}\mathcal{L}_{ext}$ is our extrapolated gradient which can be controlled by the extrapolation factor $\beta$ depending the the level of bias in the data.
+1(b): Visualization of batch construction.
+$B_{b}$ refers to a biased batch where the majority of images from class $y=1$ (top row) have attribute $a=1$ (yellow), and the majority of images from class $y=2$ (bottom row) have attribute $a=2$ (light-blue).
+$B_{lb}$ represents a less biased batch, with a more balanced attribute distribution within each class, controlled by $c$ (here $c=\frac{1}{2}$).
+$B_{rs}$ depicts a batch with group-balanced distribution and refers to the batch used in the Resampling method [16].
+$B_{ext}$ simulates GERNE’s batch with $c\cdot(\beta+1)>1$, where the dataset’s minority groups appear as majorities in the batch.
+1(c): A simplified 2D representation of gradient extrapolation where $\theta\in\mathbb{R}^{2}$.
+$\nabla_{\theta}\mathcal{L}_{b}$ is the gradient computed on $B_{b}$; training with this gradient is equivalent to training with the ERM objective.
+$\nabla_{\theta}\mathcal{L}_{lb}$ represents the gradient computed on $B_{lb}$.
+$\nabla_{\theta}\mathcal{L}_{rs}$ is the gradient computed on $B_{rs}$, which is equivalent, in expectation, to an extrapolated gradient with $c\cdot(\beta+1)=1$.
+Finally, $\nabla_{\theta}\mathcal{L}_{ext}$ is our extrapolated gradient, with the extrapolation factor $\beta$ modulating the degree of debiasing based on the strength of spurious correlations present in the dataset.
 2 Related Work
-Debiasing according to attributes annotations availability.
+Debiasing according to attribute annotations availability.
 Numerous studies have leveraged attribute annotations to mitigate spurious correlations and learning debiased representation [39, 55, 3, 51, 53].
-For instance, Group DRO [39] optimizes model performance on the worst-case group by minimizing worst-group error during training.
-While these methods are effective, obtaining attribute annotations for each sample can be extremely time-consuming and labor-intensive.
-Consequently, recent works have explored approaches that rely on limited attribute-labeled data to reduce the dependency on full annotations [18, 42, 32].
-For example, DFR [18] enhances robustness by using a small, group-balanced validation set with attribute labels to retrain the final layer of a pre-trained model.
-For cases where attribute information is only available for model selection and hyperparameter tuning [6, 16, 28, 31, 53], usually an initial model is used to separate samples based on the alignment between the label and spurious attributes.
-Samples for which this model incurs relatively low loss are considered ”easy” examples, where we expect the alignment to hold and the samples to closely resemble the majority group.
-In contrast, samples with high loss are considered ”hard” examples, and these samples tend to resemble the minority group [48].
-This process effectively creates ”easy” and ”hard” pseudo-attributes within each class, allowing debiasing methods that traditionally rely on attribute information to be applied.
-For example, JTT [28] first trains a standard ERM model and then trains a second model by upweighting the misclassified training examples (hard examples) detected by the first model.
+For instance, Group DRO [39] optimizes model performance on the worst-case group by directly minimizing worst-group loss during training.
+While effective, such methods rely on complete attribute annotations, which are often costly and labor-intensive to obtain.
+Consequently, recent works have explored approaches that reduce reliance on full annotations by using limited attribute information [18, 42, 32].
+For example, DFR [18] enhances robustness by using a small, group-balanced validation set with attribute information to retrain the final layer of a pre-trained model.
+In cases where attribute information is only available for model selection and hyperparameter tuning [6, 16, 28, 31, 54], an initial or auxiliary ERM-trained model is often used to infer the attributes by partitioning the training data into majority and minority groups.
+Samples on which the model incurs relatively low loss (i.e., high-confidence predictions) are treated as “easy” examples—where spurious correlations are likely to hold— and these examples form the majority group.
+Conversely, high-loss samples are considered “hard” examples, and typically form the minority group where such correlations may not apply [49].
+This process effectively creates “easy” and “hard” pseudo-attributes within each class, allowing debiasing methods that traditionally rely on attribute information to be applied.
+For example, JTT [28] first trains a standard ERM model and then trains a second model by upweighting the misclassified training examples detected by the first model.
 Finally, a more realistic and challenging scenario arises when attribute information is entirely unavailable [4, 43]—not accessible for training, model selection, or hyperparameter tuning—requiring models to generalize without explicit guidance on non-causal features [50].
@@
 Resampling underrepresented groups to ensure a more balanced distribution of samples [16, 19] or modifying the loss function to adjust for imbalances [38] are common examples of these techniques.
-In this work, we demonstrate in Sec.
-5.3 that although balancing techniques are effective, their performance is constrained in the presence of spurious correlations.
+We demonstrate in Sec.
+5.3 that although the balancing techniques are effective, their performance is constrained in the presence of spurious correlations.
 In contrast, our proposed debiasing approach mitigates the negative effects of spurious correlations by guiding the learning process in a debiasing direction, proving to be more effective.
 3 Problem Setup
-We consider a regular multi-class classification problem with $K$ classes and $A$ spurious attributes.
-Each input sample $x_{i}\in\mathcal{X}=\{x_{i}\mid i=1,\dots,N\}$ is associated with a class label $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$ and an attribute $a_{i}\in\mathcal{A}=\{1,\dots,A\}$, where $N$ is the total number of input samples in the dataset.
-We define a group $\mathcal{X}_{y,a}$ for $(y,a)\in\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ as the set of input samples $x_{i}$ with a class label $y$ and an attribute $a$, resulting in $|\mathcal{G}|=K\cdot A$ groups.
-We define $\mathcal{X}_{y}=\cup_{a\in\mathcal{A}}\mathcal{X}_{y,a}$ as the set of input samples $x_{i}$ with a class label $y$.
-We assume that all groups are non-empty (i.e.
-$\forall(y,a)\in\mathcal{G},\mathcal{X}_{y,a}\neq\emptyset$).
-Our goal is to learn the intrinsic features that define the labels rather than spurious features present in a biased dataset.
-This ensures robust generalization when spurious correlations are absent in the test distribution.
+We consider a standard multi-class classification problem with $K$ classes and $A$ attributes.
+Each input sample $x_{i}\in\mathcal{X}=\{x_{j}\mid j=1,\dots,N\}$ is associated with a class label $y_{i}\in\mathcal{Y}=\{1,\dots,K\}$ and an attribute $a_{i}\in\mathcal{A}=\{1,\dots,A\}$, where $N$ is the total number of samples in the dataset.
+We define a group $\mathcal{X}_{y,a}$ for $(y,a)\in\mathcal{G}=\mathcal{Y}\times\mathcal{A}$ as the set of input samples $x_{i}$ with class label $y$ and attribute $a$, resulting in $|\mathcal{G}|=K\cdot A$ groups.
+For each class $y$, we denote by $\mathcal{X}_{y}=\bigcup_{a\in\mathcal{A}}\mathcal{X}_{y,a}$ the set of all samples with label $y$.
+We assume all groups are non-empty, i.e., $\forall(y,a)\in\mathcal{G},\mathcal{X}_{y,a}\neq\emptyset$, and denote the cardinality of any group $\mathcal{X}_{m}$ by $|\mathcal{X}_{m}|$.
+Our goal is to learn the intrinsic features that define the labels, rather than spurious features present in a biased dataset, where spurious correlations are prevalent.
+This would ensure robust generalization when spurious correlations are absent in the test distribution.
 Following  [39], we aim to learn a function parameterized by a neural network $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to minimize the risk for the worst-case group:
-$f^{*}=\arg\min_{f}\max_{g\in\mathcal{G}}\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right].$ (1)
+$f^{*}=\arg\min_{f}\max_{g\in\mathcal{G}}\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right],$ (1)
 where $\ell(y,f(x))\rightarrow\mathbb{R}$ is the loss function.
 4 The Proposed Method: GERNE
+We build GERNE with the goal of mitigating the impact of spurious correlations.
 The core idea of GERNE is to sample two batches with different amounts of spurious correlations, hereafter named the biased batch $\boldsymbol{B_{b}}$ and the less biased batch $\boldsymbol{B_{lb}}$ (Fig.
@@
 Let $\mathcal{L}_{b},\mathcal{L}_{lb}$ be the losses calculated on $\boldsymbol{B_{b}}$ and $\boldsymbol{B_{lb}}$, respectively.
-We assume that utilizing the extrapolation of the gradients of these two losses towards the gradient of $\mathcal{L}_{lb}$ guides the model toward debiasing as illustrated in Fig.
+We assume that extrapolating the gradients of these two losses towards the gradient of $\mathcal{L}_{lb}$ guides the model toward debiasing as illustrated in Fig.
 1(c).
-We first present our method GERNE for training with known attributes and then generalize GERNE to the unknown attribute case.
+We first present GERNE for training with known attributes and then generalize GERNE to the unknown attribute case.
 4.1 GERNE for the Known Attributes Case
-In the following, we denote the joint distribution of $(y,a)$ in a sampled batch from the dataset as $p(y,a)$.
-We define two types of batches with different conditional attribute distributions $p(a|y)$: the biased and the less biased batches.
-Then, we define our target loss as the linear extrapolation of the two losses computed on the biased and less biased batches.
-A simplified illustration of our approach is shown in Fig.
+In the following, we denote by $p(y,a)$ the joint distribution of class label $y$ and attribute $a$ in a sampled batch.
+During training, we construct two types of batches with different conditional attribute distributions $p(a|y)$: the biased and the less biased batches.
+Our method defines the target loss as a linear extrapolation between the losses computed on these two batches.
+A simplified illustration is shown in Fig.
 1.
-Finally, we derive the link between the extrapolation factor and the risk of the worst-case group in Eq.
-1, and we theoretically define the upper and lower bounds of this factor.
+Finally, we derive the link between the extrapolation factor and the risk for the worst-case group in Eq.
+1, and theoretically define the upper and lower bounds of this factor.
 4.1.1 Sampling the biased and the less biased batches
-The biased batch and less biased batches are sampled to satisfy the following two conditions:
+The biased batch and the less biased batches are sampled to satisfy the following two conditions:
   1. 1.
-    Uniform sampling from classes, i.e., $\forall y\in\mathcal{Y},p(y)=\frac{1}{K}$
+    Uniform sampling from classes, i.e., $\forall y\in\mathcal{Y},p(y)=\frac{1}{K}.$
   2. 2.
@@
 $\alpha_{ya}=\frac{|\mathcal{X}_{y,a}|}{|\mathcal{X}_{y}|}.$ (2)
-Here $|\mathcal{X}_{y,a}|,|\mathcal{X}_{y}|$ denote the cardinality of $\mathcal{X}_{y,a},\mathcal{X}_{y}$, respectively.
 Note that to sample a biased batch, no access to the attributes is required, and uniformly sampling from $\mathcal{X}_{y}$ for each label $y$ satisfies Eq.
 2.
-While the less biased batch ($\boldsymbol{B_{lb}}$) is sampled with a conditional attribute distribution, denoted as $p_{lb}(a|y)$, which is defined to be less biased compared to the biased batch.
-That is: $\forall(y,a)\in\mathcal{G}$,
+The less biased batch ($\boldsymbol{B_{lb}}$) is sampled with a conditional attribute distribution, denoted as $p_{lb}(a|y)$, which satisfies the following: $\forall(y,a)\in\mathcal{G}$:
 $\ \min(\frac{1}{A},p_{b}(a|y))\leq p_{lb}(a|y)\leq\max(\frac{1}{A},p_{b}(a|y)).$ (3)
-Here $\mathcal{L}_{lb}$ quantifies the loss when the representation of the spurious features is partially reduced. Choosing
+That is, $B_{lb}$ exhibits a more balanced group distribution than $B_{b}$, and $\mathcal{L}_{lb}$ quantifies the loss when spurious correlations are reduced in the sampled batch. Choosing
 $p_{lb}(a|y)=(1-c)\cdot p_{b}(a|y)+c\cdot\frac{1}{A}=\alpha_{ya}+c\cdot(\frac{1}{A}-\alpha_{ya})$ (4)
@@
 We define our target loss $\mathcal{L}_{ext}$ as follows:
-$\mathcal{L}_{ext}=\mathcal{L}_{lb}+\beta(\mathcal{L}_{lb}-\mathcal{L}_{b}).$ (5)
+$\mathcal{L}_{ext}=\mathcal{L}_{lb}+\beta\cdot(\mathcal{L}_{lb}-\mathcal{L}_{b}),$ (5)
 where $\beta$ is a hyperparameter, and the loss form given the joint distribution $p(x,y,a)$ is defined as:
 $\mathcal{L}=\mathbb{E}_{(x,y,a)\sim p(x,y,a)}\left[\ell(y,f(x))\right].$ (6)
-Given the set of parameters $\theta$ of our model $f$, the gradient vector of the loss $\mathcal{L}_{ext}$ with respect to $\theta$ can be derived from Eq. 5:
-$\nabla_{\theta}\mathcal{L}_{ext}=\nabla_{\theta}\mathcal{L}_{lb}+\beta\left(\nabla_{\theta}\mathcal{L}_{lb}-\nabla_{\theta}\mathcal{L}_{b}\right).$ (7)
+Given the set of parameters $\theta$ of our model $f$, the gradient of $\mathcal{L}_{ext}$ with respect to $\theta$ can be derived from Eq. 5:
+$\nabla_{\theta}\mathcal{L}_{ext}=\nabla_{\theta}\mathcal{L}_{lb}+\beta\cdot\left(\nabla_{\theta}\mathcal{L}_{lb}-\nabla_{\theta}\mathcal{L}_{b}\right).$ (7)
 Our target gradient vector $\nabla_{\theta}\mathcal{L}_{ext}$ in Eq.
-7 is the linear extrapolation of the two gradient vectors $\nabla_{\theta}\mathcal{L}_{lb}$ and $\nabla_{\theta}\mathcal{L}_{b}$, and accordingly, we call $\beta$ the extrapolation factor.
-Since the less biased batch has a less skewed conditional attribute distribution in comparison to the biased batch according to Eq.
-3, the extrapolation of their gradients and toward the less biased gradient forms a new gradient ($\mathcal{L}_{ext}$) that leads to learning even more debiasing representation for some values of the extrapolation factor $\beta>0$.
-The concept of extraploation is shown in Fig.
+7 is a linear extrapolation of the two gradient vectors $\nabla_{\theta}\mathcal{L}_{lb}$ and $\nabla_{\theta}\mathcal{L}_{b}$, and accordingly, we refer to $\beta$ as the extrapolation factor.
+Because the less biased batch has a less skewed conditional attribute distribution compared to the biased batch (as shown in Eq.
+3), extrapolating their gradients and toward the less biased gradient forms a new gradient ($\mathcal{L}_{ext}$) that leads to learning even more debiased representation for some values of the extrapolation factor $\beta>0$.
+A visual representation of extrapolation is shown in Fig.
 1(c).
-4.1.3 GERNE as general framework for debiasing
+4.1.3 GERNE as a general framework for debiasing
 Minimizing our target loss $\mathcal{L}_{ext}$ simulates minimizing the loss of class-balanced batches with the following conditional distribution of $(y,a)\in\mathcal{G}$ :
 $p_{ext}(a|y)=\alpha_{ya}+c\cdot(\beta+1)\cdot\left(\frac{1}{A}-\alpha_{ya}\right).$ (8)
-The proof is in Appendix A.
-Based on Eq. 8, we can establish the link between GERNE and other methods for different values of $\beta$ and $c$ as illustrated in Fig. 1:
-    For $\beta=-1$, $\mathcal{L}_{ext}=\mathcal{L}_{b}$ and GERNE is equivalent to balanced-classes ERM approach.
-    For $c=1,\beta=0$, $p_{ext}(a|y)=\frac{1}{A}$; GERNE is equivalent to the resampling method [16] (upsampling the minority group).
-    For $c\cdot(\beta+1)=1$, $p_{ext}(a|y)=\frac{1}{A}$; In this case, $\mathcal{L}_{ext}$ is equivalent, in expectation, to the loss of the resampling method ($\mathcal{L}_{rs}$ in Fig.
-1).
-However, the variances of the losses differ.
-In fact, GERNE permits controlling the variance of its loss, thanks to its hyperparameters, which might help escaping sharp minima [1] and generalize better [23].
-The derivation of the variance of GERNE’s loss is detailed in Appendix C.
+We provide the full proof in Appendix A.
+Based on Eq. 8, we can establish the link between GERNE and other methods for different values of $\beta,c$:
+    For $\beta=-1$, $\mathcal{L}_{ext}=\mathcal{L}_{b}$ and GERNE is equivalent to class-balanced ERM method.
+    For $c=1$ and $\beta=0$, $p_{ext}(a|y)=\frac{1}{A}$, and GERNE matches Resampling  [16], which samples equally from all groups ($B_{rs}$ in Fig.
+1(b), with gradient of the loss computed on it denoted as $\nabla_{\theta}\mathcal{L}_{rs}$ in Fig.
+1(c)).
+    For $c\cdot(\beta+1)=1$, we also have $p_{ext}(a|y)=\frac{1}{A}$, and $\mathcal{L}_{ext}$ is, in expectation, equivalent to $\mathcal{L}_{rs}$.
+However, their loss variances differ.
+In fact, GERNE permits controlling the variance of its loss through its hyperparameters ($c,\beta$), which may help escape sharp minima  [1] and improve generalization  [23].
+The derivation of the variance of GERNE’s loss is detailed in Appendix B.
     For $c\cdot(\beta+1)>1$, $p_{ext}(a|y)>\frac{1}{A}\text{ if }\alpha_{ya}<\frac{1}{A}$ (also $p_{ext}(a|y)<\frac{1}{A}\text{ if }\alpha_{ya}>\frac{1}{A}$).
-In this case, GERNE simulates batches where the underrepresented groups (i.e.
-$\alpha_{ya}<\frac{1}{A}$) in the dataset to be oversampled in these simulated batches.
+In this case, GERNE simulates batches where the underrepresented groups (i.e., those with $\alpha_{ya}<\frac{1}{A}$) are oversampled.
 4.1.4 Upper and lower bounds of $\beta$
-Having $p_{ext}(a|y)$ in Eq. 8 in $[0,1]$, $\beta$ should satisfy:
-$\begin{split}\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\min(i_{ya}^{1},i_{ya}^{2})\leq\beta\leq\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})\end{split}$ (9)
-For: $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$.
-These limits are used when tuning $\beta,c$.
-4.1.5 Tuning $\beta$ to minimize the risk of worst-case group
-Eq. 5 can be rewritten as follows:
-$\mathcal{L}_{ext}=\frac{1}{K}\sum_{g=(y,a)\in\mathcal{G}}p_{ext}(a|y)(\beta)\cdot L_{g}$ (10)
+Having $p_{ext}(a|y)$ in Eq. 8 within $[0,1]$, $\beta$ should satisfy:
+$\begin{split}\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\min(i_{ya}^{1},i_{ya}^{2})\leq\beta\leq\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2}),\end{split}$ (9)
+where: $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$.
+These bounds are used when tuning $\beta$.
+In Appendix C, we simplify these bounds to $\beta_{\min},\beta_{\max}]=[-1,i_{y^{\prime\prime}a^{\prime\prime}}^{1}]$, where $(y^{\prime\prime},a^{\prime\prime})=\arg\max_{(y,a)\in\mathcal{G}}\alpha_{ya}$.
+Note that $\beta$ doesn’t affect $p_{ext}(a|y)$ for $\alpha_{ya}=\frac{1}{A}$ according to Eq.
+8.
+4.1.5 Tuning $\beta$ to minimize the risk for worst-case group
+Eq. 5 can be rewritten as follows (detailed in Appendix A):
+$\mathcal{L}_{ext}=\frac{1}{K}\cdot\sum_{g=(y,a)\in\mathcal{G}}p_{ext}(a|y)(\beta)\cdot L_{g},$ (10)
 where
@@
 Taking this into consideration, we define $g^{\prime}=(y^{\prime},a^{\prime})=\arg\min_{(y,a)\in\mathcal{G}}\alpha_{ya}$.
-Since $L_{g^{\prime}}$ is multiplied by $p_{ext}(a^{\prime}|y^{\prime})$ defined in Eq.
-8, increasing $\beta$ assigns more weight to $L_{g^{\prime}}$ in Eq.
-10 than $L_{g}$ with $g\neq g^{\prime}$, thereby encouraging a reduction in $L_{g^{\prime}}$ during training.
+Since $L_{g^{\prime}}$ is weighted by $p_{ext}(a^{\prime}|y^{\prime})$, increasing $\beta$ beyond $\frac{1}{c}-1$ assigns more weight to $L_{g^{\prime}}$ in Eq.
+10 than any other group loss (all groups’ losses are equally weighted when $c\cdot(\beta+1)=1$).
+This increase in $\beta$ encourages the model to prioritize reducing the loss of the underrepresented group $g^{\prime}$ during training, therefore minimizing the risk for the worst-case group.
 We outline the detailed steps of our approach for the known attribute case in Algorithm 1.
 Algorithm 1 GERNE for the known attribute case
-Input: $\mathcal{X}_{y,a}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$ and $a\in\mathcal{A}$, $f$ with initial $\theta=\theta_{0}$, ##\## epochs $E$, batch size per label $B$, ##\## classes $K$, ##\## attributes $A$, learning rate $\eta$.
-1:Choose $c\in(0,1]$, $\beta$ satisfying Eq. 9. (REC: start with $c=\frac{1}{2},\beta=1$).
-2:for epoch $=1$ to $E$ do
+Input: $\mathcal{X}_{y,a}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$ and $a\in\mathcal{A}$, $f$ with initial $\theta=\theta_{0}$, #\# epochs $E$, batch size per label $B$, #\# classes $K$, #\# attributes $A$, learning rate $\eta$.
+1:Choose $c\in(0,1]$ and $\beta\in[\beta_{\min},\beta_{\max}]$ via grid search.
+2:for epoch =1=1 to $E$ do
 3:  Biased Batch $B_{b}=\emptyset$, Less Biased Batch $B_{lb}=\emptyset$
@@
 6:   $B_{b}=B_{b}\cup B_{b}^{y,a}$
-7:         Sample a mini-batch $B_{lb}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $((1-c)~{}\alpha_{y,a}+\frac{c}{A})\cdot B$
+7:   Sample a mini-batch $B_{lb}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $((1-c)\cdot\alpha_{y,a}+\frac{c}{A})\cdot B$
 8:   $B_{lb}=B_{lb}\cup B_{lb}^{y,a}$
@@
 11:  Compute $\nabla_{\theta}\mathcal{L}_{ext}$ using Eq. 7.
-12:     Update parameters $\theta$: $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{ext}$
+12:  Update parameters (SGD): $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{ext}$
 13:end for
 4.2 GERNE for the Unknown Attributes Case
-If the attributes are not known during training, it is not possible to directly sample biased and less biased batches.
-This is because the conditional attribute distribution, $p(a|y)$, is unknown.
-Hence, we follow the previous work [31, 28, 53] by training a standard ERM model $\tilde{f}$ and using its predictions to create pseudo-attributes $\tilde{a}$.
-By training $\tilde{f}$ using the biased batches, the model learns to rely on spurious correlations, leading to biased predictions.
-Leveraging these predictions, we can classify samples into hard examples (instances with low confidence) and easy examples (instances with high confidence).
-After training $\tilde{f}$, we select a threshold $t\in(0,1)$ and derive pseudo-attributes based on model predictions as follows: For each class $y$, we compute the predictions $\tilde{y}_{i}=p(y|x_{i})=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$.
-Then, we split the predictions into two non-empty subsets.
-The first subset contains the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ predictions, and the corresponding samples of these predictions represent group $\mathcal{X}_{y,\tilde{a}=1}$.
+If the attributes are unavailable during training, it is not possible to directly sample less biased batches.
+To address this, we follow the previous work  [31, 28, 54] by training a standard ERM model $\tilde{f}$ and using its predictions to create pseudo-attributes $\tilde{a}$.
+Since $\tilde{f}$ is trained on biased batches, it tends to rely on spurious correlations, resulting in biased predictions.
+Leveraging these predictions, we classify samples into easy—those with high-confidence predictions, where the spurious correlations likely hold— and hard — those with low-confidence predictions, where the spurious correlations may not hold.
+After training $\tilde{f}$, we select a threshold $t\in(0,1)$ and construct pseudo-attributes based on model predictions as follows: For each class $y$, we compute the predictions $\tilde{y}_{i}=p(y|x_{i})=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$.
+We then split them into two non-empty subsets: The first subset contains the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ values, and the corresponding samples form the group $\mathcal{X}_{y,\tilde{a}=1}$.
 The remaining samples forms the group $\mathcal{X}_{y,\tilde{a}=2}$.
 This process ensures that each set $\mathcal{X}_{y}$ is divided into two disjoint and non-empty groups.
-Consequently, the pseudo-attribute space consists of two values, denoted as $\mathcal{\tilde{A}}=\{1,2\}$ ($\tilde{A}=2$) with $\tilde{\mathcal{G}}=\mathcal{Y}\times\mathcal{\tilde{A}}$ replacing $\mathcal{G}$ in the unknown attribute case.
-$t$ is a hyperparameter and we outline the detailed steps for the unknown case in Algorithm 2 in Appendix D.
+Consequently, the pseudo-attribute space consists of two values, denoted as $\mathcal{\tilde{A}}=\{1,2\}$ (i.e., $\tilde{A}=2$) with $\tilde{\mathcal{G}}=\mathcal{Y}\times\mathcal{\tilde{A}}$ replacing $\mathcal{G}$ in the unknown attribute case.
+$t$ is a hyperparameter, and we outline the detailed steps of GERNE for the unknown case in Appendix D.
 4.2.1 Tuning $\beta$ to control the unknown conditional distribution of an attribute $a$ in class $y$
-After creating the pseudo-attributes and defining the pseudo-groups, suppose we form a new batch of size $B$ with $\gamma\cdot B$ samples uniformly sampled form group $(y,\tilde{a}=1)$ and $(1-\gamma)\cdot B$ samples from group $(y,\tilde{a}=2)$, for $\gamma\in[0,1]$.
-The conditioal attribute distribution of an attribute $a$ given $y$ in the new formed batch is given by:
-$p_{B}(a|y)$ $=\sum_{\tilde{a}\in\mathcal{\tilde{A}}}p_{B}(\tilde{a}|y)\cdot p(a|\tilde{a},y)$ (12)
-Because the max/min value of a linear program must occur at a vertex, we have:
-$\forall\gamma\in[0,1],\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\leq p_{B}(a|y)\leq\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ (13)
-This means that if: $\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)<\frac{1}{A}(\text{or }\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)>\frac{1}{A})$, then there is no value for $\gamma$ that allows creating a batch with $p_{B}(a|y)>\frac{1}{A}$ or $p_{B}(a|y)<\frac{1}{A})$ via sampling from the pseudo-groups.
+After creating the pseudo-attributes and defining the pseudo-groups, we consider forming a new batch of size $B$ by uniformly sampling $\gamma\cdot B$ examples from group $\mathcal{X}_{y,\tilde{a}=1}$ and $(1-\gamma)\cdot B$ examples from group $\mathcal{X}_{y,\tilde{a}=2}$, where $\gamma\in[0,1],\gamma\cdot B\in\mathbb{N}$.
+The resulting conditional distribution of an attribute $a$ given $y$ in the constructed batch is:
+$p_{B}(a|y)$ $=\sum_{\tilde{a}\in\mathcal{\tilde{A}}}p_{B}(\tilde{a}|y)\cdot p(a|\tilde{a},y).$ (12)
+Because the max/min value of a linear program must occur at a vertex, we have for $p(a|\tilde{a},y)=p_{\tilde{a},y}(a)$:
+$\forall\gamma\in[0,1],\min_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)\leq p_{B}(a|y)\leq\max_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a).$ (13)
+This means that if: $\max_{\tilde{a}}p_{\tilde{a},y}(a)<\frac{1}{A}$$(\min_{\tilde{a}}p_{\tilde{a},y}(a)>\frac{1}{A})$, then there is no value for $\gamma$ can yield a batch with $p_{B}(a|y)>\frac{1}{A}$$(p_{B}(a|y)<\frac{1}{A})$ via sampling from the pseudo-groups.
 Proposition 1.
-Creating a biased and less biased batch with the pseudo-attributes $\mathcal{\tilde{A}}=2$, and with $c,\beta$ as hyperparameter, we can simulate creating batches with more controllable conditional attribute distribution (i.e.
-$p_{B}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{B}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$).
-The proof is in Appendix E.
+In case of unknown attributes, GERNE can simulate creating batches with more controllable conditional attribute distribution (i.e., $p_{B}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)\text{ or }p_{B}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)$).
+We provide the proof of this proposition in Appendix E.
 5 Experiments
-The performance of our proposed approach, GERNE, is evaluated across five computer vision datasets and one natural language processing dataset, specifically: Colored MNIST (C-MNIST) [3, 27], Corrupted CIFAR-10 (C-CIFAR-10)[15, 31], Biased FFHQ (bFFHQ) [22, 27], Waterbird [45], CelebA [30], and CivilComments [5].
-Based on the evaluation metric and experimental setup used, we divide these datasets into two groups: Datasets-1 and Datasets-2.
-Datasets-1 contains the first three datasets listed above, with the goal of assessing the performance of GERNE without any data augmentation methods.
-While Datasets-2 contains the remaining three datasets, with our implementation unified with [49].
-In the following, we first describe our experimental setup and results for each group of datasets, and then we present an ablation study for hyperparameter tuning.
+To evaluate the general applicability of GERNE, we assess its performance across five computer vision and one natural language processing benchmarks: Colored MNIST (C-MNIST)  [3, 27], Corrupted CIFAR-10 (C-CIFAR-10) [15, 31], Biased FFHQ (bFFHQ)  [22, 27], Waterbird  [46], CelebA  [30], and CivilComments  [5].
+We categorize these datasets into two groups: Datasets-1 and Datasets-2.
+Datasets-1 comprises the first three datasets mentioned above and is used to evaluate GERNE’s performance without data augmentation.
+Datasets-2 consists of the remaining three datasets, for which we follow the experimental setup described in  [50] to ensure a fair comparison.
 5.1 Experiments on Datasets-1
 Datasets.
-The C-MNIST dataset represents an extension of the MNIST dataset [26], incorporating colored digits.
-Each digit is highly correlated with a specific color, which constitutes its majority group.
-In the C-CIFAR-10 dataset, each category of images is corrupted with a specific type of texture noise [15].
-The bFFHQ dataset comprises human face images, with ”age” and ”gender” as the target and spurious attributes, respectively.
-The majority of images depicting females are labeled as ”young” while the majority of images depicting males are labeled as ”old”.
+C-MNIST is an extension of the MNIST dataset  [26] where each digit class is predominantly associated with a specific color.
+This introduces a spurious correlation between the digit label (target) and color (attribute).
+C-CIFAR-10 modifies CIFAR-10 by applying specific texture patterns to each object class  [15], making texture a spurious feature.
+Both C-MNIST and C-CIFAR-10 include versions with varying degrees of spurious correlation, reflected by the minority group ratios of 0.5%, 1%, 2%, and 5% in the training and validation sets.
+The bFFHQ dataset comprises human face images, with “age” and “gender” as the target and spurious attributes, respectively.
+The majority of female faces are young, while the majority of males are old.
+The minority group ratio in the training set is 0.5%.
 Evaluation metrics.
-For both C-MNIST and C-CIFAR-10, we train with varying ratios of the minority to majority examples (0.5%, 1%, 2%, and 5%), and we follow the evaluation setup of [31, 29] using GBA on the test set as our evaluation metric.
-For bFFHQ, we train models with a 0.5% minority ratio and evaluate the performance based on the accuracy of the minority group in line with [27].
+We follow the evaluation protocols of prior work [31, 29, 27].
+For C-MNIST and C-CIFAR-10, we report Group-Balanced Accuracy (GBA) on the test set.
+For bFFHQ, we evaluate performance based on the accuracy of the minority group.
 Baselines.
-We consider six baseline methods: For the known attribute case, we compare GERNE with Group DRO [39].
-For the unknown attribute case, our baselines are ERM [44], JTT [28], LfF [31], DFA [27], and LC [29].
+For the known attribute case, we compare GERNE with Group DRO  [39] and Resampling  [16].
+For the unknown attribute case, we consider ERM  [45], JTT  [28], LfF  [31], DFA  [27], LC  [29], and DeNetDM  [44].
 Implementation details.
-We adopt the same model architectures as the baseline methods and utilize the SGD optimizer with a momentum of 0.9 and weight decay of 0.01 across all three datasets.
-Additional implementation details in Sec.
-F.1.
+We adopt the same model architectures as the baselines and use SGD optimizer across all three datasets. More details are provided in Sec. F.1.
+Table 1: Performance comparison of GERNE and baselines on the C-MNIST, C-CIFAR-10, and bFFHQ datasets.
+We report GBA (%) with standard deviation over three trials for C-MNIST and C-CIFAR-10 across varying minority ratios, and minority group accuracy (%) for bFFHQ.
+DeNetDM results are from [44], and Resampling results are generated using GERNE with $c=1,\beta=0$.
+All other results are from  [29].
+$\checkmark/\times$ indicates known/unknown training attributes.
+The best results are marked in bold, and the second-best are underlined.
+[…]
 Results.
 Tab.
-1 reports our results against baseline methods for the known and unknown attribute cases.
-The results of baseline methods are adopted from [29].
+1 compares GERNE with baselines for both known and unknown attribute cases.
+All baseline results are adopted from  [29], except DeNetDM, which is sourced from  [44].
 When the attributes are known, GERNE outperforms Group DRO by a significant margin on C-MNIST and C-CIFAR-10 datasets.
-The improvement in performance is ranging from about 5% on C-CIFAR-10 with 5% of minority group and up to 16% on C-MNIST with 1% of minority group.
-Furthermore, we show that we outperform the resampling method [16] ($c=1,\beta=0$) for all ratios.
-More discussion added to Appendix B.
-For bFFHQ, GERNE results in an improvement of over 7% in comparison with the resampling method.
-For the unknown attribute case, our method outperforms all the baseline methods except for C-MNIST with 5% of minority group where GERNE achieves the second-best accuracy, with only a 0.18% difference from the best result, while maintaining a lower standard deviation, despite not employing any data augmentation techniques.
-Table 1: Performance comparison of GERNE and baseline methods on the C-MNIST, C-CIFAR-10, and bFFHQ datasets.
-We report the GBA (%) and standard deviation over three trials on the test set for C-MNIST and C-CIFAR-10, with varying ratios (%) of minority samples.
-For bFFHQ, we report the minority group accuracy (%).
-Baseline results are sourced from [29] as the same experimental settings are adopted.
-/$\checkmark$/$\times$ / × indicate the presence/absence of attribute information in the training set, respectively.
-The best results are marked in bold, and the second-best are underlined.
-Methods Group Info C-MNIST C-CIFAR-10 bFFHQ 0.5 1 2 5 0.5 1 2 5 Group DRO 63.12 68.78 76.30 84.20 33.44 38.30 45.81 57.32 - GERNE (ours with $c=1,\beta=0$) 77.68 $\pm$0.89 84.36 $\pm$0.21 88.15 $\pm$0.11 91.98 $\pm$0.08 45.10 $\pm$0.60 50.08 $\pm$0.42 54.85 $\pm$0.30 62.16 $\pm$0.05 69.80 $\pm$1.33 GERNE (ours) 77.79 $\pm$0.90 84.47 $\pm$0.37 88.30 $\pm$0.20 92.16 $\pm$0.10 45.34 $\pm$0.60 50.84 $\pm$0.17 55.51 $\pm$0.10 62.40 $\pm$0.27 77.10 $\pm$0.90 ERM $\times$ 35.19 $\pm$3.49 52.09 $\pm$ 2.88 65.86 $\pm$ 3.59 82.17 $\pm$ 0.74 23.08 $\pm$ 1.25 28.52 $\pm$ 0.33 30.06 $\pm$ 0.71 39.42 $\pm$ 0.64 56.70 $\pm$ 2.70 JTT $\times$ 53.03 $\pm$ 3.89 62.90 $\pm$ 3.01 74.23 $\pm$ 3.21 84.03 $\pm$ 1.10 24.73 $\pm$ 0.60 26.90 $\pm$ 0.31 33.40 $\pm$ 1.06 42.20 $\pm$ 0.31 65.30 $\pm$ 2.50 LfF $\times$ 52.50 $\pm$ 2.43 61.89 $\pm$ 4.97 71.03 $\pm$ 1.14 84.79 $\pm$ 1.09 28.57 $\pm$ 1.30 33.07 $\pm$ 0.77 39.91 $\pm$ 1.30 50.27 $\pm$ 1.56 62.20 $\pm$ 1.60 DFA $\times$ 65.22 $\pm$ 4.41 81.73 $\pm$ 2.34 84.79 $\pm$ 0.95 89.66 $\pm$ 1.09 29.75 $\pm$ 0.71 36.49 $\pm$ 1.79 41.78 $\pm$ 2.29 51.13 $\pm$ 1.28 63.90 $\pm$ 0.30 LC $\times$ 71.25 $\pm$ 3.17 82.25 $\pm$ 2.11 86.21 $\pm$ 1.02 91.16 $\pm$ 0.97 34.56 $\pm$ 0.69 37.34 $\pm$ 1.26 47.81 $\pm$ 2.00 54.55 $\pm$ 1.26 69.67 $\pm$ 1.40 GERNE (ours) $\times$ 77.25 $\pm$ 0.17 83.98 $\pm$ 0.26 87.41 $\pm$ 0.31 90.98 $\pm$ 0.13 39.90$\pm$ 0.48 45.60$\pm$ 0.23 50.19$\pm$ 0.18 56.53 $\pm$ 0.32 70.87 $\pm$ 0.61
+The improvement in performance ranges from about 5% on C-CIFAR-10 with 5% of minority group and up to 16% on C-MNIST with 1% of minority group.
+Furthermore, GERNE outperforms Resampling  [16] by over 13% on bFFHQ and consistently surpasses it across all other ratios.
+Our explanation behind GERNE’s superior performance over Resampling is that the latter tends to present the majority and minority groups equally in the sampled batches during training, and the model $f$ tends to prioritize learning the easy-to-learn spurious features associated with the majority group (e.g., the colors in C-MNIST), leading to learning biased representation and poorer generalization.
+In contrast, GERNE undermines learning the spurious features by directing the learning process more in the debiasing direction, thanks to the extrapolation factor.
+For the unknown attribute case, GERNE outperforms all baselines, except on C‑MNIST with 5% of minority group (ranks second), while maintaining a lower standard deviation.
+At this 5% minority ratio, LC achieves slightly higher accuracy—likely benefiting from its use of data augmentation to increase the diversity of the samples in the minority group.
+We exclude DeNetDM’s results on C-MNIST, as the authors use a different version of this dataset.
 5.2 Experiments on Datasets-2
 Datasets.
-We evaluate GERNE on three commonly used datasets[49]: Waterbirds [45], CelebA [30] and CivilComments [5].
+Waterbirds  [46] contains bird images with spurious correlations between bird type and background: Most waterbirds appear with water backgrounds, while most landbirds appear with land backgrounds.
+CelebA  [30] involves classifying hair color (blond, non-blond), with gender (male, female) as the spurious attribute: Most blond images depict females.
+CivilComments  [5] is a binary toxic comment classification dataset, where the spurious attribute marks references to eight different demographic identities (male, female, LGBTQ, Christian, Muslim, other religions, Black, and White).
 Evaluation metrics.
-We follow the same evaluation strategy in [49] for model selection and hyperparameter tuning: When attributes are known in both training and validation, we use the worst-group accuracy of the test set.
-When attributes are unknown in training, but known in validation, we use the worst-group accuracy of the validation set.
-When attributes are unknown in both training and validation, we use the worst-class accuracy of the validation set.
+We follow the same evaluation strategy from  [50] for model selection and hyperparameter tuning.
+When attributes are known in both training and validation, we use the worst-group test accuracy as the evaluation metric.
+When attributes are unknown in training, but known in validation, we use the worst-group validation accuracy.
+When attributes are unavailable in both, we use the worst-class validation accuracy.
 Baselines.
-For each dataset, we select the three best performing methods reported in [49].
+For each dataset, we select the best three methods reported in  [50].
 We end up with ERM [45], Group DRO [39], DFR  [18], LISA  [51], ReSample  [19], Mixup  [53], ReWeightCRT  [21], ReWeight  [19], CBLoss  [7], BSoftmax [36] and SqrtReWeight  [50].
@@
 Implementation details.
-We employ the same data augmentation techniques, optimizers and pretrained models described in [49]. Further details can be found in Sec. F.2.
+We employ the same data augmentation techniques, optimizers and pretrained models described in  [50]. Further details are in Sec. F.2.
 Results
 Tab.
-2 shows the worst-group accuracy of the test set for GERNE against the baseline methods under the evaluation strategy explained above.
-In the case of known attributes, GERNE achieves the highest performance on the CelebA and CivilComments datasets and ranks second on Waterbirds, following DFR.
-In case of unknown attributes in training set but known in validation, our approach again achieves the best results on the Waterbirds and CivilComments datasets and remains competitive on CelebA, closely following the top two baseline results.
-Notably, DFR uses the validation set to train the model, whereas GERNE only uses it for model selection and hyperparameter tuning.
-We include a comparison of our method’s performance against DFR, where GERNE also uses the validation set for training, in Appendix G.
-In the scenario where attributes are unknown in both the training and validation sets, our approach achieves the best results on the Waterbirds and CelebA datasets.
-However, we observe a significant drop in accuracy on CelebA compared to the second case (unknwon attributes in training but known in validation), while this drop is less pronounced on Waterbirds.
-This can be explained by the worst-class accuracy evaluation metric.
-In the validation set of CelebA, the majority examples in class 1 exhibit spurious correlations, leading to selection process to favor the majority group while disregarding the minority group.
-However, the validation set of Waterbirds has balanced groups within each class, resulting in only a slight performance drop for GERNE between the second and third case.
-This highlights the critical role of having access to the attributes in the validation set or having a group-balanced validation set for model selection and hyperparameter tuning when aiming for better results with GERNE.
-Table 2: Performance comparison of GERNE and baseline methods on the Waterbirds, CelebA, and CivilComments datasets.
-We report the worst-group test accuracy (%) and standard deviation over three trials on the test of each dataset.
+2 shows the worst-group accuracy (WGA) of the test set for GERNE compared to the baseline methods under the evaluation strategy explained above.
+In the known attributes case, GERNE achieves the highest accuracy on CelebA and CivilComments, and ranks second on Waterbirds, following DFR.
+In case of unknown attributes in the training set but known in validation, our approach again attains the best results on Waterbirds and CivilComments datasets and remains competitive on CelebA, closely following the top two baselines’ results.
+In particular, DFR uses the validation set to train the model, whereas GERNE employs it only for model selection and hyperparameter tuning.
+We include a comparison between DFR and GERNE when using the validation set for training in Appendix G.
+When attributes are unknown in both the training and validation sets, GERNE achieves the best results on Waterbirds and CelebA.
+However, we observe a significant drop in accuracy on CelebA compared to the second case (known attributes only in validation), while this drop is less pronounced on Waterbirds.
+The difference can be attributed to the use of worst-class accuracy as the evaluation metric.
+In CelebA’s validation set, the majority of blond hair images exhibit spurious correlations (female images), leading the model selection process to favor the majority group while disregarding the minority group.
+In contrast, the validation set in Waterbirds is group-balanced within each class, leading to only a slight decrease in performance between the second and third case.
+This highlights the critical role of having access to the attributes in the validation set —or at least a group-balanced validation set— for model selection when using GERNE.
+Table 2: Performance comparison of GERNE and baseline methods on Waterbirds, CelebA, and CivilComments.
+We report the worst-group test accuracy (%) and standard deviation over three trials for each dataset.
 Baseline results are sourced from  [50] as the same experimental settings are adopted.
-$\checkmark$/$\checkmark$ / ✓ indicates known attributes in training and validation sets.
-×$\times$/$\checkmark$ / ✓ in validation set only, and ×/$\times$/$\times$ / × in neither.
+$\checkmark/\checkmark$ denotes known attributes in training and validation sets.
+$\times/\checkmark$ indicates attributes are known only in the validation set, while $\times/\times$ signifies that attributes are unknown in both sets.
 Best results are highlighted in bold, and the second-best are underlined.
-Methods Group Info Waterbirds CelebA CivilComments train attr./val attr.
-ERM ✓/✓ 69.10 $\pm$ 4.70 62.60 $\pm$ 1.50 63.70 $\pm$ 1.50 Group DRO ✓/✓ 78.60 $\pm$ 1.00 89.00 $\pm$ 0.70 70.60 $\pm$ 1.20 ReWeight ✓/✓ 86.90 $\pm$ 0.70 89.70 $\pm$ 0.20 65.30 $\pm$ 2.50 ReSample ✓/✓ 77.70 $\pm$ 1.20 87.40 $\pm$ 0.80 73.30 $\pm$ 0.50 CBLoss ✓/✓ 86.20 $\pm$ 0.30 89.40 $\pm$ 0.70 73.30 $\pm$ 0.20 DFR ✓/✓ 91.00 $\pm$ 0.30 90.40 $\pm$ 0.10 69.60 $\pm$ 0.20 LISA ✓/✓ 88.70 $\pm$ 0.60 86.50 $\pm$ 1.20 73.70 $\pm$ 0.30 GERNE (ours) ✓/✓ 90.20 $\pm$ 0.22 91.98 $\pm$ 0.15 74.65 $\pm$ 0.20 ERM $\times$/✓ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/✓ 73.10 $\pm$ 0.40 78.50 $\pm$ 1.10 69.50 $\pm$ 0.70 ReWeight $\times$/✓ 72.50 $\pm$ 0.30 81.50 $\pm$ 0.90 69.90 $\pm$ 0.60 DFR $\times$/✓ 89.00 $\pm$ 0.20 86.30 $\pm$ 0.30 63.90 $\pm$ 0.30 Mixup $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 LISA $\times$/✓ 78.20 $\pm$ 0.40 57.80 $\pm$ 0.80 66.10 $\pm$ 1.30 BSoftmax $\times$/✓ 74.10 $\pm$ 0.90 83.30 $\pm$ 0.30 69.40 $\pm$ 1.20 ReSample $\times$/✓ 70.00 $\pm$ 1.00 82.20 $\pm$ 1.20 68.20 $\pm$ 0.70 CnC $\times$/✓ 88.50$\pm$ 0.30 88.80$\pm$ 0.90 68.90$\pm$ 2.10 GERNE (ours) $\times$/✓ 90.21 $\pm$ 0.42 86.28 $\pm$ 0.12 71.00 $\pm$ 0.33 ERM $\times$/$\times$ 69.10 $\pm$ 4.70 57.60 $\pm$ 0.80 63.20 $\pm$ 1.20 Group DRO $\times$/$\times$ 73.10 $\pm$ 0.40 68.30 $\pm$ 0.90 61.50 $\pm$ 1.80 DFR $\times$/$\times$ 89.00 $\pm$ 0.20 73.70 $\pm$ 0.80 64.40 $\pm$ 0.10 Mixup $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 LISA $\times$/$\times$ 77.50 $\pm$ 0.70 57.80 $\pm$ 0.80 65.80 $\pm$ 1.50 ReSample $\times$/$\times$ 70.00 $\pm$ 1.00 74.10 $\pm$ 2.20 61.00 $\pm$ 0.60 ReWeightCRT $\times$/$\times$ 76.30 $\pm$ 0.20 70.70 $\pm$ 0.60 64.70 $\pm$ 0.20 SqrtReWeight $\times$/$\times$ 71.00 $\pm$ 1.40 66.90 $\pm$ 2.20 68.60 $\pm$ 1.10 CRT $\times$/$\times$ 76.30 $\pm$ 0.80 69.60 $\pm$ 0.70 67.80 $\pm$ 0.30 GERNE (ours) $\times$/$\times$ 89.88 $\pm$ 0.67 74.24 $\pm$ 2.51 63.10 $\pm$ 0.22
+[…]
 5.3 GERNE vs. Balancing Techniques
 Balancing techniques have been shown to achieve state-of-the-art results, while remaining easy to implement  [16, 50].
-Notably, resampling methods often outperform reweighting strategies when combined with stochastic gradient algorithms [2].
-Our results, as presented in Tab.
-1, demonstrate that GERNE outperforms resampling (GERNE with $c=1,\beta=0$) when the evaluation metric is Group-Balanced Accuracy or Accuracy on minority group.
-More discussion in Appendix B.
+While Resampling often outperforms Reweighting when combined with stochastic gradient algorithms [2], we show in Tab.
+1 that GERNE consistently outperforms Resampling in both group-balanced accuracy (GBA) and minority group accuracy.
 This highlights the flexibility of GERNE to adapt to maximize both metrics, and its superior performance in comparison to Resampling and other balancing techniques, as further supported by the results in Tab.
 2.
-In Appendix C, we provide a detailed ablation study that compares GERNE to an equivalent (in term of loss expectation)’sampling + weighting’ approach, and show how GERNE can leverage its controllable loss variance (tuned by the hyperparameters $c,\beta$) to escape sharp minima.
+In Appendix B, we provide a detailed ablation study comparing GERNE to an equivalent “sampling+weighting” approach with matching loss expectation, and demonstrate how GERNE can leverage its controllable loss variance (by the hyperparameters $c,\beta$) to escape sharp minima.
 5.4 Ablation Study
@@
 The value of $\beta$ in Eq.
-7 has a significant impact on the performance of our method in debiasing the model (i.e., leading the training process in a debiased direction and avoid learning spurious features).
+7 plays a critical role in guiding the model toward learning debiased representation (i.e., reducing reliance on spurious features and improving generalization).
 In Fig.
-2, we show how tuning $\beta$ affects the learning process in the case of the C-MNIST dataset with $0.5\%$ of minority group in the known attribute case.
+2, we illustrate the effect of tuning $\beta$ on the learning process using C-MNIST with $0.5\%$ of minority group in the known attributes case.
 We show results for $\beta\in\{-1,0,1,1.2\}\text{ with }c=0.5$.
 For $\beta=-1$, our target loss $\mathcal{L}_{ext}$ in Eq.
-5 equals the biased loss $\mathcal{L}_{b}$, which leads to learning a biased model that exhibits high accuracy in the majority group, yet demonstrates poor performance on both the minority group and the unbiased test set.
-As $\beta$ increases (e.g.
-$\beta=0,\beta=1$), the model starts learning more intrinsic features.
+5 equals the biased loss $\mathcal{L}_{b}$, which leads to learning a biased model that exhibits high accuracy on the majority group, yet demonstrates poor performance on both the minority group and the unbiased test set.
+As $\beta$ increases (e.g., $\beta=0,\beta=1$), the model starts learning more intrinsic features.
 This is evident from the improved performance on the minority group in the validation set, as well as on the unbiased test set.
-However, as the extrapolation factor $\beta$ continues to increase, the model begins to exhibit higher variance during the training process as shown for $\beta=1.2$, ultimately leading to divergence when $\beta$ exceeds the upper bound defined in Eq.
-9 which is equal to $1.22$ in this case.
-By comparing the accuracies of Minority/Majority training groups in case $\beta=0,\beta=1$, we can see that both cases have around $100\%$ accuracy on minority but higher accuarcy on majority for $\beta=0$.
-However, we notice a better generalization when $\beta=1$.
-This highlights the importance of directing the training process to the right direction early in training while overfitting is expected as well.
-Figure 2: The impact of tuning $\beta$ in debiasing the model $\beta\in\{-1,0,1,1.2\}$ on debiasing the model.
+However, as the extrapolation factor $\beta$ continues to increase, the model begins to exhibit higher variance during the training process, as shown for $\beta=1.2$, ultimately leading to divergence when $\beta$ exceeds the upper bound defined in Eq.
+9 ($1.22$ in this case).
+While GERNE appears to be sensitive to small variations in $\beta$ (e.g., 1.2 to 1.22), we show in Appendix C that $\beta_{\max}$ is inversely proportional to $c$ and $A$.
+This implies that decreasing $c$ allows for a wider feasible range of $\beta$.
+By comparing the accuracies on minority and majority training groups in case $\beta=0,\beta=1$, we can see that both cases have around $100\%$ accuracy on minority but higher accuracy on majority for $\beta=0$.
+However, $\beta=1$ results in better generalization overall.
+This highlights the importance of directing the training process toward a debiased direction early in training, especially when overfitting is likely to occur on the minority group (e.g., when it contains very few samples).
+Figure 2: The impact of tuning $\beta\in\{-1,0,1,1.2\}$ on debiasing the model.
 On the left column, we plot the training losses $\mathcal{L}_{b},\mathcal{L}_{lb}$ and the target loss $\mathcal{L}_{ext}$.
-On the right column, we plot the average accuracy of the minority and majority groups in both training and validation, as well as the average accuracy of the unbiased test set.
+On the right column, we plot the average accuracy of the minority and majority groups in both training and validation sets, as well as the average accuracy of the unbiased test set.
 Each plot represents the mean and standard deviation calculated over three runs with different random seeds.
@@
 We first train a biased model $\tilde{f}$, and use its predictions to generate the pseudo-attributes for five different values of the threshold $t$.
-Let’s refer to the pseudo-groups with $\tilde{a}=1$ as the pseudo-minority group.
-For each threshold, we tune $\beta$ to achieve the best average accuracy on test set.
-Simultaneously, we compute the average precision and recall for the minority group, as defined in Eq.
-14.
+Let’s refer to the pseudo-groups with $\tilde{a}=1$ as the pseudo-minority groups.
+For each threshold, we tune $\beta$ to achieve the best average test accuracy.
+Simultaneously, we compute the average precision and recall for the minority group.
 As shown in Fig.
-3, with $t=0.0005$, the average precision reaches 1, indicating that all the samples in the pseudo-minority group are from the minority group.
-However, these samples constitute less than 20% of the total number of samples in the minority group, as indicated by the average recall.
+3, with $t=5\times 10^{-4}$, the average precision reaches 1, indicating that all the samples in the pseudo-minority groups are from the minority groups.
+However, these samples constitute less than 20% of the total minority groups, as indicated by the average recall.
 Despite this, GERNE achieves a high accuracy of approximately 70%, remaining competitive with other methods reported in Tab.
-1 while using only a very limited number of minority samples ($t=$0.0005 corresponds to about 28 samples versus 249 minority samples out of 55,000 samples in the training set).
-As $t$ increases to 0.001 and 0.003, precision remains close to 1 while increasing the number of minority samples in the pseudo-minority group.
-This increase introduces more diversity among minority samples within the pseudo-minority group, allowing for lower $\beta$ values to achieve the best average accuracy on test set.
-However, for even higher thresholds, such as $t=$0.01, minority samples constitute less than 40% in the pseudo-minority group, prompting a need to revert to higher $\beta$ values.
-We conclude that identifying the minority group is of utmost importance for achieving optimal results (high average precision and high recall) and this agrees with the results presented in both Tab.
+1 while using only a very limited number of minority samples ($t=5\times 10^{-4}$ corresponds to about 28 samples versus 249 minority samples out of 55,000 samples in the training set).
+As $t$ increases to $10^{-3}$ and $3\times 10^{-3}$, precision remains close to 1 while increasing the number of minority samples in the pseudo-minority groups.
+This increase introduces more diversity among minority samples within the pseudo-minority groups, allowing for lower $\beta$ values to achieve the best average test accuracy.
+However, for even higher thresholds, such as $t=10^{-2}$, minority samples constitute less than 40% in the pseudo-minority groups, prompting a need to revert to higher $\beta$ values.
+We conclude that identifying the samples of minority groups (high average precision and high recall) is of utmost importance for achieving optimal results, and this agrees with the results presented in both Tab.
 1, Tab.
-2 where we achieve the best results in the case of known attributes.
-Figure 3: The effect of choosing the threshold $t$ to generate the pseudo attributes on $\beta$ for best performance.
-Average Precision $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ total in }(y,\tilde{a}=1)},$ (14)
-Average Recall $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\frac{\#\text{ minority in }(y,\tilde{a}=1)}{\#\text{ minority in }(y)}$
+2, where we achieve the best results in the known attributes case.
+Figure 3: The effect of the threshold $t$ used to generate pseudo-attributes on the extrapolation factor $\beta$ and model performance.
+We plot the average precision and recall over pseudo-minority groups $(y,\tilde{a}=1)$, averaged across all classes $y$.
+For each $(y,\tilde{a}=1)$, precision is defined as the fraction of minority samples among all samples in that group, and recall is the fraction of those minority samples relative to all minority samples in class $y$.
+We also report the best achievable test accuracy, along with the corresponding extrapolation factor $\beta$, across different threshold values.
 6 Conclusion
-In this work, we introduce GERNE, a novel debiasing approach designed to mitigate spurious correlations.
-By sampling two types of batches with varying amounts of spurious correlations, we define a debiasing direction using the gradients of the losses computed on these two batches.
-The extrapolated gradient vector, controlled by the extrapolation factor, is then used to update the model parameters.
-We demonstrate that GERNE can be adapted to optimize the Group-Balanced Accuracy (GBA) metric or the Worst-Group Accuracy (WGA) metric, leading to improved generalization and superior performance compared to state-of-the-art methods on both known and unknown attribute cases, without relying on data augmentation techniques, on multiple benchmarks.
-Notably, GERNE provides a general framework that encompasses common methods such as ERM and resampling, making it a valid approach for training models even on unbiased datasets.
-Furthermore, GERNE allows for control over the degree of loss variance during training, increasing the likelihood of escaping sharp minima and consequently improving generalization.
-Although effective, GERNE remains limited in its utilization of the diversity within the majority group, where spurious correlations are present.
-This raises a new research question: Can we identify a more effective debiasing direction based on the two previously mentioned gradients?
-Future work will aim to address this question.
-Additionally, we will explore strategies to dynamically update the extrapolation factor during training to reduce reliance on hyperparameter tuning, and investigate novel methods to better estimate the attributes in the unknown attribute case.
+We introduce GERNE, a novel debiasing approach that effectively mitigates spurious correlations by leveraging an extrapolated gradient update.
+By defining a debiasing direction from loss gradients computed on batches with varying degrees of spurious correlations, GERNE’s tunable extrapolation factor allows optimizing either Group-Balanced Accuracy (GBA) or Worst-Group Accuracy (WGA).
+Our comprehensive evaluations across vision and NLP benchmarks demonstrate GERNE’s superior performance over state-of-the-art methods, both for known and unknown attribute cases, without data augmentation.
+Furthermore, GERNE offers a general framework that encompasses methods like ERM and Resampling, extending its applicability to unbiased datasets.
+Future work will explore dynamic adaptation of the extrapolation factor and refine attribute estimation for the unknown attributes case.
+Acknowledgments.
+This work was funded by the Carl Zeiss Foundation within the project Sensorized Surgery, Germany (P2022-06-004).
+Maha Shadaydeh is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Individual Research Grant SH 1682/1-1.
 Supplementary Material
-Appendix A Conditional attribute distribution of the extrapolated loss: Proof of Eq. 8
+Appendix A The Extrapolated Conditional Attribute Distribution of GERNE:
 $\mathcal{L}$ in Eq. 6 can be written as:
-[…]
+$\mathcal{L}=\int\ell(y,f(x))p(y)\,p(a|y)\,p(x|a,y)\,dx\,dy\,da.$
 Therefore, $\mathcal{L}_{b},\mathcal{L}_{lb}$ can be written as:
-[…]
-[…]
-Where $p_{b}(a|y),p_{lb}(a|y)$ defined in Eq. 2, Eq. 4, respectively. Substituting Eq. 15, Eq. 16 in Eq. 5:
+$\mathcal{L}_{b}=\int\ell(y,f(x))p(y)\,p_{b}(a|y)\,p(x|a,y)\,dx\,dy\,da,$ (14)
+$\mathcal{L}_{lb}=\int\ell(y,f(x))p(y)\,p_{lb}(a|y)\,p(x|a,y)\,dx\,dy\,da.$ (15)
+Where $p_{b}(a|y),p_{lb}(a|y)$ defined in Eq.
+2, Eq.
+4, respectively, and $x$ is uniformly sampled within each group $\mathcal{X}_{y,a}$ (i.e., $p(x|y,a)=\frac{1}{|\mathcal{X}_{y,a}|}$).
+Substituting Eq.
+14, Eq.
+15 in Eq.
+5:
 $\mathcal{L}_{ext}$ $=\int\ell(y,f(x))\,p(y)\,\Big{(}p_{lb}(a|y)$
@@
 $\quad p(x|a,y)\,dx\,dy\,da$
-[…]
+$=\int\ell(y,f(x))\,p(y)\,p_{ext}(a|y)\cdot p(x|a,y)\,dx\,dy\,da.$
 Then:
 $p_{ext}(a|y)=\alpha_{ya}+c\cdot(\beta+1)\cdot(\frac{1}{A}-\alpha_{ya}).$
-Appendix B GERNE versus the resampling method [16]
-In Tab.
-1, GERNE achieves higher GBA compared to the special case of GERNE with $c=1,\beta=0$ (resampling method) for both C-MNIST and C-CIFAR-10 datasets.
-Our explanation behind GERNE superior performance is that resampling method tend to present the majority and minority groups equally in the batch, and the model $f$ tends to prioritize learning the easy-to-learn spurious features associated with the majority group.
-For instance, the color in C-MNIST.
-While GERNE undermines learning the spurious correlations by directing the learning process more in the debiasing direction thanks to the extrapolation factor.
-Appendix C GERNE versus an equivalent sampling and weighting approach
-We compare GERNE with an equivalent (in term of loss expectation) sampling+weighting method, which we refer to as ’SW’. For simplicity, we assume the following:
+Furthermore, we can write $\mathcal{L}_{ext}$ as follows:
+$\mathcal{L}_{\text{ext}}$ $=\mathbb{E}_{y\sim p(y)}\left[\mathbb{E}_{a\sim p_{\text{ext}}(a|y)}\left[\mathbb{E}_{x\sim p(x|a,y)}\left[\ell(y,f(x))\right]\right]\right]$
+$=\frac{1}{K}\cdot\sum_{g=(y,a)\in\mathcal{G}}p_{\text{ext}}(a|y;\beta)\cdot L_{g},$
+with $L_{g}=\mathbb{E}_{x\sim p(x|(y,a)=g)}\left[\ell(y,f(x))\right]$ by using the discrete expectations over $y$ and $a|y$, with $p(y)=\frac{1}{K}$.
+Appendix B GERNE Versus an Equivalent Sampling and Weighting Approach
+We compare GERNE with an equivalent (in terms of loss expectation) sampling+weighting method, which we refer to as “SW”. For simplicity, we assume the following:
   1. 1.
-    A binary classification task where the number of classes equals the number of attributes (i.e. $K=A=2$).
+    A binary classification task where the number of classes equals the number of attributes (i.e., $K=A=2$).
   2. 2.
-    The attributes are known, and the classes are balanced. (i.e. $|\mathcal{X}_{y=1}|=|\mathcal{X}_{y=2}|$).
+    The attributes are known, and the classes are balanced (i.e., $|\mathcal{X}_{y=1}|=|\mathcal{X}_{y=2}|$).
   3. 3.
-    The majority of samples which hold the spurious correlation in each class are aligned with the class label (i.e. $|\mathcal{X}_{y,a=y}|>|\mathcal{X}_{y,a\neq y}|$).
+    The majority of samples that hold the spurious correlation in each class are aligned with the class label (i.e., $|\mathcal{X}_{y,a=y}|>|\mathcal{X}_{y,a\neq y}|$), and the dataset is highly biased (i.e., $\frac{|\mathcal{X}_{y,a\neq y}|}{|\mathcal{X}_{y,a=y}|}\ll 1$).
   4. 4.
-    The dataset is highly biased. In other words, $\frac{|\mathcal{X}_{y,a\neq y}|}{|\mathcal{X}_{y,a=y}|}\ll 1$.
-  5. 5.
-    In a highly biased dataset, best performance is coupled with overpresenting the minority (conflicting samples according to assumption 1.) in early stages of training.
-Therefore, an overfitting on the minority is expected before the overfitting on the majority.
-We refer to the expected loss of the majority (aligned) samples as $\mathcal{L}_{A}$, and the expected loss of the minority (conflicting) as $\mathcal{L}_{C}$.
+    In a highly biased dataset, best performance is coupled with overrepresenting the minority groups (as illustrated in Fig.
+2) in early stages of training.
+Therefore, an overfitting on the minority groups is expected before the overfitting on the majority groups.
+We refer to the expected loss of the majority samples as $\mathcal{L}_{A}$, and the expected loss of the minority as $\mathcal{L}_{C}$.
 For GERNE, we sample two batches: biased and less biased batch, each of size $B$.
@@
 5, $\mathcal{L}_{ext}$ can be written as:
-$\mathcal{L}_{ext}=(1+\beta)\cdot\mathcal{L}_{lb}-\beta\cdot\mathcal{L}_{b}$ (17)
-Since the biased batch reflects the inherent bias present in the dataset, under the third assumption, we can approximate $\mathcal{L}_{b}$ by $\mathcal{L}_{A}$, neglecting the loss on the very few conflicting samples in the batch.
+$\mathcal{L}_{ext}=(1+\beta)\cdot\mathcal{L}_{lb}-\beta\cdot\mathcal{L}_{b}.$ (16)
+Since the biased batch reflects the inherent bias present in the dataset, under the third assumption, we can approximate $\mathcal{L}_{b}$ by $\mathcal{L}_{A}$, neglecting the loss on the very few samples of the minority group in the batch.
 Therefore, we have:
-$\mathcal{L}_{b}\approx\mathcal{L}_{A}$ (18)
+$\mathcal{L}_{b}\approx\mathcal{L}_{A}.$ (17)
 Following the third assumption and the conditional attribute distribution in Eq.
-4, we can approximate the composition of the less biased batch as follows: a proportion of $(1-\frac{c}{2})$ of the samples in the less biased batch are drawn from the aligned samples, while a proportion of $\frac{c}{2}$ of the samples from the minority group.
+4, we can approximate the composition of the less biased batch as follows: a proportion of $(1-\frac{c}{2})$ of the samples in the less biased batch are drawn from the aligned samples (majority group), while a proportion of $\frac{c}{2}$ of the samples from the minority group.
 This leads to the following approximation:
-$\mathcal{L}_{lb}\approx(1-\frac{c}{2})\cdot\mathcal{L}_{A}+\frac{c}{2}\cdot\mathcal{L}_{C}$ (19)
+$\mathcal{L}_{lb}\approx(1-\frac{c}{2})\cdot\mathcal{L}_{A}+\frac{c}{2}\cdot\mathcal{L}_{C}.$ (18)
 Substituting Eq. 17, Eq. 18 into Eq. 16:
 $\mathcal{L}_{ext}\approx\frac{2-c\cdot(1+\beta)}{2}\cdot\mathcal{L}_{A}+\frac{c\cdot(1+\beta)}{2}\cdot\mathcal{L}_{C}.$ (19)
-We consider the following ’SW’ approach:
-  • Sampling step : we sample an ’SW’ batch of size $B$ similar to the less biased batch in GERNE.
-Where $(1-\frac{c}{2})$ of the batch samples are from the majority group (aligned samples) and $\frac{c}{2}$ from minority (conflicting samples).
+We consider the following “SW” approach:
+  • Sampling step: we sample an “SW” batch of size $B$, analogous to the less biased batch in GERNE, where $(1-\frac{c}{2})$ of the samples are from the majority group (aligned samples) and $\frac{c}{2}$ from minority group.
   • Weighting step: we compute the loss $\mathcal{L}_{sw}$ over the sampled batch as follows:
-    $\mathcal{L}_{sw}=w\cdot\mathcal{L}_{A}+(1-w)\cdot\mathcal{L}_{C},w=\frac{2-c\cdot(1+\beta)}{2}$ (21)
-    where $\mathcal{L}_{A}$ is computed over aligned samples in the ’SW’ batch and $\mathcal{L}_{C}$ is computed over the conflicting samples.
-Let’s compute the variance of the two losses:
+    $\mathcal{L}_{sw}=w\cdot\mathcal{L}_{A}+(1-w)\cdot\mathcal{L}_{C},w=\frac{2-c\cdot(1+\beta)}{2},$ (20)
+    where $\mathcal{L}_{A}$ is computed over the samples of the majority group in the “SW” batch and $\mathcal{L}_{C}$ is computed over the samples of the minority group.
+Both $\mathcal{L}_{ext}$ and $\mathcal{L}_{sw}$ in Eq. 19, Eq. 20 are equivalent in expectation. Let’s compute the variance of the two losses:
 $Var(\mathcal{L}_{sw})$ $=w^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})+(1-w)^{2}\cdot Var(\mathcal{L}_{C}^{c/2})$
-$\quad+2\cdot w\cdot(1-w)\cdot Cov(\mathcal{L}_{A}^{1-c/2},\mathcal{L}_{C}^{c/2})$ (22)
-where $Var(\mathcal{L}^{m})$ means the variance computed over $m\cdot B$ samples where $B$ is the batch size. For simplicity, we refer to $Var(\mathcal{L}^{1})\text{ as }Var(\mathcal{L})$
-Following the fourth assumption, when the model overfits on the conflicting samples (i.e.
-$\mathcal{L}_{C}\approx 0$), we can approximate both $Var(\mathcal{L}_{C}),Cov(\mathcal{L}_{A},\mathcal{L}_{C})$ to zero.
-Therefore:
+$\quad+2\cdot w\cdot(1-w)\cdot Cov(\mathcal{L}_{A}^{1-c/2},\mathcal{L}_{C}^{c/2}),$ (21)
+[…]
 $Var(\mathcal{L}_{sw})$ $\approx w^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=\frac{w^{2}}{1-\frac{c}{2}}\cdot Var(\mathcal{L}_{A})=$
-$\quad(\frac{2-c\cdot(1+\beta)}{2})^{2}\cdot\frac{2}{2-c}\cdot Var(\mathcal{L}_{A})$ (23)
+$\quad(\frac{2-c\cdot(1+\beta)}{2})^{2}\cdot\frac{2}{2-c}\cdot Var(\mathcal{L}_{A}).$ (22)
 From Eq. 16:
@@
 $\quad-2\cdot(1+\beta)\cdot\beta\cdot Cov(\mathcal{L}_{lb},\mathcal{L}_{b})$
-$\quad\geq((1+\beta)\cdot\sqrt{Var(\mathcal{L}_{lb})}-\beta\cdot\sqrt{Var(\mathcal{L}_{b})})^{2}$ (24)
-Note that the the inequality reduces to an equality in Appendix C if $Cov(\mathcal{L}_{lb},\mathcal{L}_{b})=\sqrt{Var(\mathcal{L}_{lb})}\cdot\sqrt{Var(\mathcal{L}_{b})}$.
-The covariance term ($Cov$) can be controlled by the number of shared samples between the biased and less biased batches.
-If all the aligned samples in the less biased batch are included in the sampled biased batch (i.e.
-the less biased batch is created by replacing some aligned samples by conflicting ones), we get maximum value for the $Cov$.
-From Eq. 19:
-$Var(\mathcal{L}_{lb})\approx(1-\frac{c}{2})^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=(1-\frac{c}{2})\cdot Var(\mathcal{L}_{A})$ (25)
-And from Eq. 18
-$Var(\mathcal{L}_{b})\approx Var(\mathcal{L}_{A})$ (26)
-Finally, substituting Eq. 25 and Eq. 26 in Appendix C:
-$Var(\mathcal{L}_{ext})\geq((1+\beta)\cdot\sqrt{1-\frac{c}{2}}-\beta)^{2}\cdot Var(\mathcal{L}_{A})$ (27)
+$\quad\geq((1+\beta)\cdot\sqrt{Var(\mathcal{L}_{lb})}-\beta\cdot\sqrt{Var(\mathcal{L}_{b})})^{2}.$ (23)
+Note that the inequality reduces to an equality in Appendix B if $Cov(\mathcal{L}_{lb},\mathcal{L}_{b})=\sqrt{Var(\mathcal{L}_{lb})}\cdot\sqrt{Var(\mathcal{L}_{b})}$.
+The covariance term $Cov(.,.)$ can be controlled by the number of shared samples between the biased and less biased batches.
+If all the aligned samples in the less biased batch are included in the biased batch (i.e., the less biased batch is created by replacing some samples of the majority group with samples from the minority group), we then maximize $Cov(.,.)$.
+From Eq.
+18, we can write:
+$Var(\mathcal{L}_{lb})\approx(1-\frac{c}{2})^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=(1-\frac{c}{2})\cdot Var(\mathcal{L}_{A}),$ (24)
+and from Eq. 17:
+$Var(\mathcal{L}_{b})\approx Var(\mathcal{L}_{A}).$ (25)
+Finally, substituting Eq. 24 and Eq. 25 in Appendix B:
+$Var(\mathcal{L}_{ext})\geq((1+\beta)\cdot\sqrt{1-\frac{c}{2}}-\beta)^{2}\cdot Var(\mathcal{L}_{A}).$ (26)
 According to the fourth assumption, we are interested in the range where $c\cdot(\beta+1)\geq 1$.
@@
 9, we obtain $\beta\in\left[\frac{1-c}{c},\frac{2-c}{c}\right]$.
-As $\beta\to\frac{2-c}{c}$, the representation of aligned samples simulates a vanishing representation (according to Eq.
+As $\beta\to\frac{2-c}{c}$, the representation of the aligned samples is vanishing (according to Eq.
 8) in the sampled batches, which leads to $\mathcal{L}_{A}>0$.
-Assuming a limited and non-vanishing variance $Var(\mathcal{L}_{A})$ (i.e.
-the model outputs a non-constant prediction for samples from the majority group), we have:
-$\beta\to\frac{2-c}{c}\implies Var(\mathcal{L}_{sw})\approx 0,Var(\mathcal{L}_{ext})\neq 0$ for $c\in(0,1]$.
-This non-vanishing variance of GERNE’s loss, if controlled with tuning $\beta$ to ensure stability, gives the model the chance of escape sharp minima similar to gradient random perturbation [1] and therefore, improve generalization [33, 23, 34].
-Appendix D Algorithm 2.
+Assuming a limited and non-vanishing variance $Var(\mathcal{L}_{A})$ (i.e., the model outputs non-constant predictions for samples from the majority group), we have:
+$\beta\to\frac{2-c}{c}\implies Var(\mathcal{L}_{sw})\approx 0,\text{ while }Var(\mathcal{L}_{ext})\neq 0$ for $c\in(0,1]$.
+This non-vanishing variance of GERNE’s loss, if controlled with tuning $\beta$ to ensure stability, gives the model the chance of escaping sharp minima similar to gradient random perturbation [1] and therefore, improves generalization [33, 23, 34].
+Appendix C Simplifying the Bounds of $\beta$
+We aim to simplify the upper and lower bounds of $\beta$ in Eq. 9. We start by simplifying the upper bound:
+$\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2}),$
+where $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$, and under the following constraints: $\forall y\in\mathcal{Y},\sum_{a}\alpha_{ya}=1$, $\alpha_{ya}\in]0,1[\setminus\{\frac{1}{A}\}$, $A\geq 2$, and $c\in(0,1]$.
+We note that $i_{ya}^{1}$ is a decreasing, and $i_{ya}^{2}$increasing function in $\alpha_{ya}$.
+We can show that if $\alpha_{ya}<\frac{1}{A}$, then $i_{ya}^{2}>i_{ya}^{1}$, and if $\alpha_{ya}>\frac{1}{A}$, then $i_{ya}^{1}>i_{ya}^{2}$.
+We conclude with the following:
+$\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}<\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})=i_{y^{\prime}a^{\prime}}^{2},\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}>\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})=i_{y^{\prime\prime}a^{\prime\prime}}^{1},$
+where:
+$\alpha_{y^{\prime}a^{\prime}}=\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\end{subarray}}\alpha_{ya}<\frac{1}{A},\alpha_{y^{\prime\prime}a^{\prime\prime}}=\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\end{subarray}}\alpha_{ya}>\frac{1}{A}.$
+Since $\sum_{k}\alpha_{y^{\prime}k}=1$, we have:
+$\sum_{k\neq a^{\prime}}\alpha_{y^{\prime}k}=1-\alpha_{y^{\prime}a^{\prime}},$
+which implies that there exists $j\neq a^{\prime}$ such that:
+$\alpha_{y^{\prime}j}\geq\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}.$
+Given that $\alpha_{y^{\prime}a^{\prime}}<\frac{1}{A}$ and $A\geq 2$, it follows that
+$\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}>\frac{1}{A},$
+and hence $\alpha_{y^{\prime}j}>\frac{1}{A}$. Therefore, we have
+$\max(i_{y^{\prime}j}^{1},i_{y^{\prime}j}^{2})=i_{y^{\prime}j}^{1}=-\frac{\alpha_{y^{\prime}j}}{c\cdot(\frac{1}{A}-\alpha_{y^{\prime}j})}-1.$
+Since $i_{ya}^{1}$ is a decreasing function in $\alpha_{ya}\in(\frac{1}{A},1]$, we have
+$i_{y^{\prime}j}^{1}\leq-\frac{\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}}{c\cdot(\frac{1}{A}-\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1})}-1=\frac{1-\alpha_{y^{\prime}a^{\prime}}}{c\cdot(\frac{1}{A}-\alpha_{y^{\prime}a^{\prime}})}-1=i_{y^{\prime}a^{\prime}}^{2}.$
+Thus: $i_{y^{\prime}j}^{1}\leq i_{y^{\prime}a^{\prime}}^{2}$, and since $i_{y^{\prime\prime}a^{\prime\prime}}^{1}\leq i_{y^{\prime}j}^{1}$, we conclude that the upper bound of $\beta$ is: $\beta_{\max}=i_{y^{\prime\prime}a^{\prime\prime}}^{1}$.
+For the lower bound of $\beta$, we can follow the same previous step, but we choose $\beta=-1$ as the lower bound ($\beta=-1$ satisfies Eq.
+9 and simulates ERM training as shown in Sec.
+4.1.3).
+In conclusion, for the known attributes case, we tune $\beta$ within the interval: $\beta_{\min},\beta_{\max}]=[-1,i_{y^{\prime\prime}a^{\prime\prime}}^{1}]$.
+The upper bound $\beta_{\max}$ is inversely proportional to both $c,A$.
+Consequently, larger values of $c$ reduce the feasible range for the extrapolation factor, making GERNE appear more sensitive to small variations in $\beta$.
+Appendix D Algorithm 2
 Algorithm 2 GERNE for the unknown attribute case
-Input: $\mathcal{X}_{y}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$, $f$ with initial $\theta=\theta_{0},\tilde{\theta}=\tilde{\theta_{0}}$ (parameters of the biased model $\tilde{f}$), ##\## epochs $E$, batch size per class label $B$, ##\## classes $K$, ##\## attributes $\tilde{A}=2$, learning rate $\eta$.
-1:Training $\tilde{f}$ on biased batches with class balanced accuracy CBA $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\mathbb{P}_{x|y}(y=\arg\max_{y^{\prime}\in\mathcal{Y}}\tilde{f}_{y^{\prime}}(x))\quad$ as the evaluation metric for model selection.
-2:Select a threshold $t$ and create the pseudo-groups $\tilde{\mathcal{G}}$ by following the steps in Sec. 4.2.
-3:Follow Algorithm 1. with: $\mathcal{G}\leftarrow\tilde{\mathcal{G}}$.
+Input: $\mathcal{X}_{y}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$, $f$ with initial $\theta=\theta_{0},\tilde{\theta}=\tilde{\theta_{0}}$ (parameters of the biased model $\tilde{f}$), #\# epochs $E$, batch size per class label $B$, #\# classes $K$, #\# attributes $\tilde{A}=2$, learning rate $\eta$.
+1:Training $\tilde{f}$ on biased batches with class balanced accuracy CBA $\frac{1}{K}\sum_{y\in\mathcal{Y}}\mathbb{P}_{x|y}(y=\arg\max_{y^{\prime}\in\mathcal{Y}}\tilde{f}_{y^{\prime}}(x))\quad$ as the evaluation metric for model selection.
+2:Select a threshold $t$ and create the pseudo-groups $\tilde{\mathcal{G}}$: For each class $y$, we compute the predictions $\tilde{y}_{i}=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$.
+We then form the pseudo-minority group $\mathcal{X}_{y,\tilde{a}=1}$ as the samples $x_{i}$ that have the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ predictions.
+The remaining samples form the pseudo-majority group $\mathcal{X}_{y,\tilde{a}=2}=\mathcal{X}_{y}\setminus\mathcal{X}_{y,\tilde{a}=1}$.
+3:Follow Algorithm 1 with $\tilde{\mathcal{G}}$ replacing $\mathcal{G}$.
+We consider a higher upper bound for $\beta$ than $\beta_{\max}$ derived in Appendix C as justified by the proof of Proposition 1.
+in Appendix E.
 Appendix E Proposition 1.
-Creating both biased and less biased batches using the pseudo-groups $\mathcal{\tilde{G}}$, and with $\beta$ as hyperparameter, we can simulate batches with a more controllable conditional attribute distribution.
+Creating both biased and less biased batches using the pseudo-groups $\mathcal{\tilde{G}}$, and with $\beta$ as a hyperparameter, we can simulate batches with a more controllable conditional attribute distribution.
 Specifically, for $(y,a)\in\mathcal{G}$, we can achieve scenarios where $p_{ext}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{ext}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ as opposed to Eq.
@@
 Proof.
-We define $\tilde{\alpha}_{y\tilde{a}}$ the same way as in Eq.
-2 for the created pseudo-groups: $\tilde{\alpha}_{y\tilde{a}}=\frac{|\mathcal{X}_{y,\tilde{a}}|}{|\mathcal{X}_{y}|}$.
-For a constant $c$, we create the less biased batch as in Eq.
+We define $\alpha_{y\tilde{a}}$ the same way as in Eq.
+2 for the created pseudo-groups: $\alpha_{y\tilde{a}}=\frac{|\mathcal{X}_{y,\tilde{a}}|}{|\mathcal{X}_{y}|}$.
+For a constant $c$ and $A=2$, we create the less biased batch as in Eq.
 4:
-$p_{lb}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (28)
+[…]
 Similar to Eq. 8, the conditional attribute distribution $p_{ext}(\tilde{a}|y)$ is given by:
-$p_{ext}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (29)
+[…]
 We can write $p_{ext}(a|y)$ as follows:
@@
 Placing Eq. 28 in Eq. 29, we get
-$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdot p(a|\tilde{a},y)+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdot p(a|\tilde{a},y).$ (31)
-For $p(a|\tilde{a}=1,y)\neq p(a|\tilde{a}=2,y)$ and $\tilde{\alpha}_{y1}\neq\frac{1}{2}$, to make $p_{ext}(a|y)=p$ for some $p\in[0,1]$, we can choose:
-$\beta=\frac{p-\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdotp(a|\tilde{a},y)}{\sum_{\tilde{a}\in\tilde{\mathcal{A}}}c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdotp(a|\tilde{a},y)}-1.\blacksquare$ (32)
+$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\alpha_{y\tilde{a}}\cdot p(a|\tilde{a},y)+c\cdot(\beta+1)\cdot(\frac{1}{2}-\alpha_{y\tilde{a}})\cdot p(a|\tilde{a},y).$ (30)
+For $p(a|\tilde{a}=1,y)\neq p(a|\tilde{a}=2,y)$ and $\alpha_{y1}\neq\frac{1}{2}$, to make $p_{ext}(a|y)=p$ for some $p\in[0,1]$, we tune $\beta$ until reaching the $\beta_{target}$ defined as follows:
+$\beta_{target}=\frac{p-\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\alpha_{y\tilde{a}}\cdot p(a|\tilde{a},y)}{\sum_{\tilde{a}\in\tilde{\mathcal{A}}}c\cdot(\frac{1}{2}-\alpha_{y\tilde{a}})\cdot p(a|\tilde{a},y)}-1.$ (31)
 Discussion.
-When $\tilde{\alpha}_{y1}=\frac{1}{2}$, our algorithm is equivalent to sampling uniformly from $\mathcal{X}_{y}$ and equally from classes.
-When $p(a|\tilde{a}=1,y)=p(a|\tilde{a}=2,y)$, it implies that $\tilde{f}$ has distributed the samples with attribute $a$ and class $y$ equally between the two pseudo-groups.
-However, in practice, this is precisely the scenario that $\tilde{f}$ is designed to avoid.
+When $\alpha_{y1}=\frac{1}{2}$, our algorithm is equivalent to sampling uniformly from $\mathcal{X}_{y}$ and equally from classes.
+When $p(a|\tilde{a}=1,y)=p(a|\tilde{a}=2,y)$, it implies that $\tilde{f}$ has distributed the samples with attribute $a$ and class $y$ equally between the two pseudo-groups $\mathcal{X}_{y,\tilde{a}\in\mathcal{A}}$.
+However, in practice, this is exactly the scenario that $\tilde{f}$ is designed to avoid.
 Specifically, if $a$ represents the presence of spurious attributes (i.e., the majority group), it is likely that $p(a|\tilde{a}=1,y)<p(a|\tilde{a}=2,y)$.
 Conversely, when $a$ represents the absence of spurious features (i.e., the minority group), we would expect $p(a|\tilde{a}=1,y)>p(a|\tilde{a}=2,y)$.
-In fact, $\tilde{f}$ is explicitly trained to exhibit a degree of bias, which inherently disrupts the equality above.
+In fact, $\tilde{f}$ is explicitly trained to exhibit a degree of bias, which inherently disrupts the above equality.
 Appendix F Implementation Details
-F.1 Implementation details on Datasets-1
-For the C-MNIST [3, 27], we deploy a multi-layer perceptron (MLP) with three fully connected layers, while for C-CIFAR-10 [15, 31] and bFFHQ [22, 27], we employ a pretrained ResNet-18 model [14] (pretrained on ImageNet1K [8]) as the backbone.
-The Stochastic Gradient Descent (SGD) optimizer, with a momentum of 0.9 and a weight decay of 0.01, is applied across all three datasets.
-Batch sizes are configured as follows: 100 per group/pseudo-group for C-MNIST and C-CIFAR-10, and 32 for bFFHQ.
-Learning rates are set to 0.1 for C-MNIST in the known attribute case and 0.01 in the unknown attribute case.
-For C-CIFAR-10 and bFFHQ, the learning rate is set to 0.0001.
-For GERNE in the known attribute case, we present results from two experimental configurations.
-In the first experiment, we set $c=1\text{ and }\beta=0$, which corresponds to resampling [16] from groups, i.e., training on $\mathcal{L}_{lb}$ without extrapolation.
-In the second experiment, $\beta$ is tuned for $c\in\{\frac{1}{2},1\}$.
-In the unknown attribute case, $t$ is an additional hyperparameter to be tuned.
-We avoid using any data augmentations as certain transformations can unintentionally fail to preserve the original label.
-For example, flips and rotations in C-MNIST can distort labels (e.g., a rotated ”6” appearing as a ”9”) [41].
+F.1 Implementation Details for Datasets-1
+For the C-MNIST [3, 27], we deploy a multi-layer perceptron (MLP) with three fully connected layers, while for C-CIFAR-10 [15, 31] and bFFHQ [22, 27], we employ ResNet-18 model [14], pretrained on ImageNet1K [8], as the backbone.
+We apply the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay of $10^{-2}$ across all three datasets.
+We set the batch size to 100 per group or pseudo-group for both C-MNIST and C-CIFAR-10, and to 32 for bFFHQ.
+For C-MNIST, we use a learning rate of $10^{-1}$ in the known attributes case and $10^{-2}$ in the unknown attributes case.
+For C-CIFAR-10 and bFFHQ, we use a learning rate of $10^{-4}$.
+In the unknown attribute case, we treat the threshold $t$ as an additional hyperparameter that requires tuning.
+We avoid using any data augmentations, as certain transformations can unintentionally fail to preserve the original label.
+For example, flips and rotations in C-MNIST can distort labels (e.g., a rotated “6” appearing as a “9”) [41].
 For training $\tilde{f}$ in case of unknown attributes in the training set, we employ the same model architecture as $f$, with modifications to the hyperparameters: the weight decay is doubled, and the learning rate is reduced to one-tenth of the learning rate used to train $f$.
-The loss function used is the cross-entropy loss in all the experiments.
-F.2 Implementation details on Datasets-2
-To ensure a fair comparison of GERNE with other methods in [49], we adhere to the same experimental settings.
-For the Waterbirds [45] and CelebA [30] datasets, we utilize a pretrained ResNet-50 model [14] as the backbone, while for CivilComments [5], we use a pretrained BERT model [9].
-Each backbone is followed by an MLP layer with $K$ output neurons.
-We employ SGD with a momentum of 0.9 and a weight decay of 0.01 for Waterbirds and CelebA, while for CivilComments, we use AdamW [24] optimizer with a weight decay of 0.0001 and a tunable dropout rate.
+The loss function used is the Cross-entropy loss for all the experiments.
+F.2 Implementation Details for Datasets-2
+To ensure a fair comparison between GERNE and the methods in [50], we adopt the same experimental settings.
+For Waterbirds [46] and CelebA [30] datasets, we use a pretrained ResNet-50 model [14] as the backbone, while for CivilComments [5], we use a pretrained BERT model [9].
+We append an MLP classification head with $K$ outputs.
+We employ SGD with a momentum of 0.9 and a weight decay of $10^{-2}$ for Waterbirds and CelebA.
+For CivilComments, we use AdamW [24] optimizer with a weight decay of $10^{-4}$ and a tunable dropout rate.
 We set batch sizes to 32 for both Waterbirds and CelebA and 5(16) per group(pseudo-group) for CivilComments.
-The learning rates are configured as follows: 0.0001 for Waterbirds and CelebA, and 0.00001 for CivilComments.
+The learning rates are configured as follows: $10^{-4}$ for Waterbirds and CelebA, and $10^{-5}$ for CivilComments.
 Additionally, we set the bias reduction factors $c$ to 0.5 for Waterbirds and CelebA and to 1 for CivilComments.
@@
 We employ the Cross-entropy loss as the loss function across all experiments.
-For experiments with unknown attributes in both the training and validation sets, we limit the hyperparameter $t$ search space to the interval $[0,\frac{1}{2}]$
-Appendix G Evaluating GERNE under limited attribute information
-To further demonstrate the effectiveness of GERNE in scenarios with limited access to samples with attribute information, we conduct two experiments on the CelebA dataset.
-In these experiments, we exclude the training set and only use the validation set with its attribute information for training.
-We follow the same settings and implementation details described earlier.
+For experiments with unknown attributes in both the training and validation sets, we limit the search space for $t$ to the interval $0,\frac{1}{2}]$.
+Appendix G Evaluating GERNE with Limited Attribute Information
+To further demonstrate the effectiveness of GERNE in scenarios with limited access to samples with known attributes, we conduct two experiments on the CelebA dataset using only the validation set with its attribute information for training (excluding the training set).
+We follow the same experimental setup and implementation details outlined in Sec.
+F.2.
 As part of the implementation, we first tune the hyperparameters using the designated evaluation metric.
@@
 Finally, we report the average worst-group test accuracy and standard deviation across these runs.
-Experiment 1 - Evaluation on test set.
-In this experiment, $f$ is trained using the full validation set.
-The worst-group accuracy on the test set is used as the evaluation metric.
+Evaluation on the Test Set.
+In this experiment, we train the model $f$ using the entire validation set and use the worst-group test accuracy as the evaluation metric.
 This setup represents the best possible performance achievable when relying solely on the validation set for training.
-Experiment 2 - Cross-validation.
-we divide the validation set into three non-overlapping folds, ensuring that each fold preserves the same group distribution as the original validation set (i.e., we randomly and equally distribute samples from each group across the folds).
-We use two folds to train $f$ and reserve the remaining fold for hyperparameter tuning and model selection, using the worst-group accuracy on this fold as the evaluation metric.
-We repeat this process three times, with each fold serving as the validation fold exactly once.
-We summarize the average worst-group test accuracy and standard deviation across all nine runs (three folds × three seeds) in Tab.
+Cross-Validation.
+In this experiment, we divide the validation set into three non-overlapping folds, ensuring that each fold preserves the same group distribution as the original validation set.
+Specifically, we randomly and evenly distribute samples from each group in the validation set across the folds.
+We train $f$ using two of the folds, and use the remaining fold for hyperparameter tuning and model selection, where the worst-group accuracy on this fold serves as the evaluation metric.
+We repeat this process three times so that each fold is used once as the validation fold.
+Finally, we report the average worst-group test accuracy and standard deviation across all nine runs (three folds × three seeds) in Tab.
 3.
-We also compare these results with DFR, a method that trains the last layer of the model on the validation set after performing ERM training on the training set.
-GERNE consistently achieves state-of-the-art results, demonstrating its robustness and effectiveness even with severely limited attribute information.
-Table 3: Performance Comparison of GERNE and DFR using the validation set for training.
+We compare the results of GERNE with DFR, a method that trains the final layer on the validation set following ERM training on the training set.
+GERNE consistently achieves state-of-the-art results, demonstrating its robustness and effectiveness even under limited attribute information.
+Table 3: Performance comparison of GERNE and DFR. GERNE uses only the validation set for training. We report the worst-group test accuracy (%) and standard deviation over three trials.
 Method WGA on Test Set (%)
 DFR 86.30 $\pm$ 0.30
-GERNE - Evaluation on test set 90.97 $\pm$ 0.35
-GERNE - Cross-validation 88.63 $\pm$ 0.59
+GERNE — Evaluation on Test Set 90.97 $\pm$ 0.35
+GERNE — Cross-Validation 88.63 $\pm$ 0.59